### v1

# Towards Generalisable Time Series Understanding Across Domains

Abstract
In natural language processing and computer vision, self-supervised pre-training on large datasets unlocks foundational model capabilities across domains and tasks. However, this potential has not yet been realised in time series analysis, where existing methods disregard the heterogeneous nature of time series characteristics. Time series are prevalent in many domains, including medicine, engineering, natural sciences, and finance, but their characteristics vary significantly in terms of variate count, inter-variate relationships, temporal dynamics, and sampling frequency. This inherent heterogeneity across domains prevents effective pre-training on large time series corpora. To address this issue, we introduce OTiS, an open model for general time series analysis, that has been specifically designed to handle multi-domain heterogeneity. We propose a novel pre-training paradigm including a tokeniser with learnable domain-specific signatures, a dual masking strategy to capture temporal causality, and a normalised cross-correlation loss to model long-range dependencies. Our model is pre-trained on a large corpus of $640,187$ samples and $11$ billion time points spanning $8$ distinct domains, enabling it to analyse time series from any (unseen) domain. In comprehensive experiments across $15$ diverse applications - including classification, regression, and forecasting - OTiS showcases its ability to accurately capture domain-specific data characteristics and demonstrates its competitiveness against state-of-the-art baselines. Our code and pre-trained weights are publicly available at https://github.com/oetu/otis.
1 Introduction
In natural language processing (NLP) or computer vision (CV), generalisable language features, e.g. semantics and grammar (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023), or visual features, e.g. edges and shapes (Geirhos et al., 2019; Dosovitskiy et al., 2021; Oquab et al., 2024), are learned from large-scale data. Self-supervised pre-training paradigms are designed to account for the specific properties of language (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023) or imaging (Zhou et al., 2022; Cherti et al., 2023; Oquab et al., 2024), unlocking foundational model capabilities that apply to a wide range of domains and downstream tasks. This potential, however, remains largely unrealised in time series due to the lack of self-supervised pre-training paradigms that account for the heterogeneity of time series across domains.
Time series are widespread in everyday applications and play an important role in various domains, including medicine (Pirkis et al., 2021), engineering (Gasparin et al., 2022), natural sciences (Ravuri et al., 2021), and finance (Sezer et al., 2020). They differ substantially with respect to the number of variates, inter-variate relationships, temporal dynamics, and sampling frequency (Fawaz et al., 2018; Ismail Fawaz et al., 2019; Ye & Dai, 2021; Wickstrøm et al., 2022). For instance, standard $10$-$20$ system electroencephalography (EEG) recordings come with up to $256$ variates (Jurcak et al., 2007), while most audio recordings have only $1$ (mono) or $2$ (stereo) variates. Weather data shows high periodicity, whereas financial data is exposed to long-term trends. Both domains encompass low-frequency data recorded on an hourly (278278278\,278mHz), daily ($12\,\mu$Hz), or even monthly (386386386\,386nHz) basis, while audio data is sampled at high frequencies of 44.144.144.1\,44.1kHz or more. Overall, this heterogeneity across domains renders the extraction of generalisable time series features difficult (Fawaz et al., 2018; Gupta et al., 2020; Iwana & Uchida, 2021; Ye & Dai, 2021).
While most existing self-supervised pre-training methods for time series are limited to single domains (Wu et al., 2021; 2022a; Nie et al., 2023; Dong et al., 2024; Jiang et al., 2024), recent works propose simple techniques to incorporate time series from multiple domains (Yang et al., 2024; Das et al., 2024; Woo et al., 2024; Liu et al., 2024). These works for instance crop all time series into segments of unified size (Jiang et al., 2024), resample them to a uniform frequency (Yang et al., 2024), or analyse each variate of a multi-variate time series independently (Liu et al., 2024). While these naive techniques address differences in sampling frequency and variate count, they degrade the original time series and neglect the critical inter-variate relationships and temporal dynamics required for effective real-world analysis. Consequently, there is a clear need for pre-training strategies that adequately handle heterogeneity in time series to unlock foundational model capabilities.
In this work, we propose a novel multi-domain pre-training paradigm that addresses the full spectrum of time series heterogeneity across domains. Our approach facilitates the comprehensive extraction of generalisable features from diverse time series. Pre-trained on a large corpus of publicly available data, our open model for general time series analysis (OTiS) can be fine-tuned on limited data of any (unseen) domain to perform a variety of downstream tasks, as showcased in Figure 1.
Figure 1: Overview of OTiS. Pre-trained on a large corpus of time series from diverse domains, OTiS enables general time series analysis. Its domain-specific tokeniser addresses time series heterogeneity across domains - including different numbers of variates, inter-variate relationships, temporal dynamics, and sampling frequencies - by learning unique domain signatures. After pre-training, the model can be fine-tuned on limited data from any domain, including previously unseen ones, to perform various tasks such as classification, regression, and forecasting.
Our key contributions can be summarised as follows:
  1. 1.
    We present OTiS, an open model for general time series analysis, with our entire pipeline and pre-trained weights publicly available at https://github.com/oetu/otis.
  2. 2.
    We propose a novel pre-training paradigm based on masked data modelling to address heterogeneity in multi-domain time series. Our approach includes a novel tokeniser with learnable signatures to capture domain-specific data characteristics, a dual masking strategy to learn temporal causality, and a normalised cross-correlation loss to model long-range dependencies.
  3. 3.
    We pre-train OTiS on a large corpus of publicly available time series from $8$ domains, spanning medicine, engineering, natural sciences, and finance. With $640,187$ samples and $11$ billion time points, this corpus represents diverse time series characteristics, enabling generalisable feature extraction.
  4. 4.
    We evaluate OTiS across $15$ downstream applications, including classification, regression, and forecasting. Our comprehensive analysis demonstrates that OTiS accurately captures domain-specific data characteristics and is competitive with both specialised and general state-of-the-art (SOTA) models, achieving new SOTA performance in $10$ tasks. Notably, none of the baselines is capable of performing all the tasks covered by OTiS.
2 Related Works
2.1 Self-Supervised Learning for Time Series
Time series vary significantly across domains, with differences in the number of variates, inter-variate relationships, temporal dynamics, and sampling frequencies. Due to this inherent heterogeneity, most existing works focus on pre-training models within a single domain (Oreshkin et al., 2019; Tang et al., 2020; Wu et al., 2021; Zhou et al., 2021; Wu et al., 2022a; Woo et al., 2022; Yue et al., 2022; Zhang et al., 2022; Li et al., 2023; Nie et al., 2023; Zeng et al., 2023; Dong et al., 2024). To develop more generalisable time series models, recent methods have explored multi-domain pre-training by addressing certain aspects of the heterogeneity, such as differences in variate count and sampling frequency. For instance, Liu et al. (2024) treat each variate in multi-variate time series independently to standardise generative tasks like forecasting, while Goswami et al. (2024) extend uni-variate analysis to discriminative tasks like classification. Similarly, Jiang et al. (2024) and Yang et al. (2024) standardise time series by cropping them into segments of predefined size and resampling them to a uniform frequency, respectively, to enable general classification capabilities in medical domains.
While partially addressing time series heterogeneity, these methods limit model capabilities for general time series analysis. Standardisation techniques like cropping or resampling may distort inter-variate relationships, temporal dynamics, and long-range dependencies. Additionally, many of these approaches are tailored to specific applications, such as generative tasks (Das et al., 2024; Liu et al., 2024; Woo et al., 2024), or focused on particular domains like medicine (Jiang et al., 2024; Yang et al., 2024). Moreover, recent foundational models (Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) focus on uni-variate analysis, ignoring crucial inter-variate relationships essential for real-world applications, such as disease prediction (Schoffelen & Gross, 2009; Wu et al., 2022b). Our study aims to overcome these limitations by fully addressing heterogeneity of multi-domain time series, establishing a foundation for general time series analysis across domains and tasks.
2.2 Time Series Tokenisation
Transformers (Vaswani et al., 2017) have emerged as the preferred architecture for foundational models in NLP and CV due to their scalability (Kaplan et al., 2020; Gordon et al., 2021; Alabdulmohsin et al., 2022), enabling the training of models in the magnitude of $100$ billion parameters (Chowdhery et al., 2023; Touvron et al., 2023; Oquab et al., 2024; Ravi et al., 2024). To utilise a Transformer for time series analysis, a tokeniser is required to map the time series into a compact latent space. Current methods (Jin et al., 2023; Nie et al., 2023; Zhou et al., 2023; Das et al., 2024; Goswami et al., 2024; Jiang et al., 2024; Liu et al., 2024; Woo et al., 2024; Yang et al., 2024) follow established techniques from NLP and CV, dividing time series into patches of pre-defined size. These patches are then flattened into a $1$D sequence, with positional embeddings used to retain positional information. While uni-variate models (Nie et al., 2023; Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) consider only temporal positions, multi-variate approaches (Woo et al., 2024; Yang et al., 2024; Jiang et al., 2024) account for both temporal and variate positions. However, none of these methods address the unique characteristics of variates, mistakenly assuming that the relationships between variates are identical across domains. Our work seeks to adapt the tokenisation process to preserve the domain-specific relationships between variates.
3 Methods
Figure 2: Architecture of OTiS. During pre-training, batches of time series from diverse domains are processed using a domain-specific tokeniser. This tokeniser splits a time series into fixed-size patches, which are then embedded using a patch projector shared across all variates and domains. A temporal embedding and a domain-specific variate embedding are added to each patch embedding. A dual masking strategy is employed to mask the resulting input tokens. The reconstruction of the multi-domain input tokens is guided using an auxiliary normalised cross-correlation (NCC) loss.
In this work, we present a novel multi-domain pre-training paradigm that enables generalisable feature extraction from large, heterogeneous time series corpora. We introduce a domain-specific tokeniser with learnable signatures to address heterogeneity in multi-domain time series, as described in Section 3.1. We tailor masked data modelling (MDM) for multi-domain time series to pre-train our open model for general time series analysis (OTiS) on a large, heterogeneous corpus, as detailed in Section 3.2. In particular, we propose normalised cross-correlation as a loss term to capture global temporal dynamics in time series, as explained in Section 3.3. Moreover, we introduce a dual masking strategy to capture bidirectional relationships and temporal causality, essential for general time series analysis, as described in Section 3.4. After pre-training, we fine-tune OTiS on limited data to perform a variety of downstream tasks in any - including previously unseen - domain, as outlined in Section 3.5. A graphical visualisation of our method is provided in Figure 2.
3.1 Domain-Specific Tokeniser
Overview.
Assume a time series sample ${\bm{X}}\in\mathbb{R}^{V_{S}\times T}$ from domain $S$, where $V_{S}$ denotes the number of variates specific to $S$ and $T$ denotes the number of time points. We randomly crop or zero-pad ${\bm{X}}$ to a fixed context length of $\overline{T}$ time points. We then split it into $T^{\prime}$ temporal patches of size $P$ along the time dimension, resulting in $V_{S}\cdot T^{\prime}$ patches ${\bm{x}}_{v,t}\in\mathbb{R}^{1\times P}$, where $v\in\{1,\dots,V_{S}\}$ and $t\in\{1,\dots,T^{\prime}\}$.
Next, we embed these patches using a shared patch projector across all variates and domains, resulting in patch embeddings $e^{\mathcal{P}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}\in\mathbb{R}^{1\times D}$, where $D$ denotes the model dimension. The patch projector consists of a $1$D convolutional layer followed by layer normalisation and GELU activation.
The permutation-equivariant nature of Transformers (Vaswani et al., 2017) requires the use of positional embeddings to accurately capture the inherent relationships in the input data. Initially introduced for $1$D textual token sequences (Vaswani et al., 2017), positional embeddings simply introduce an ordering into the input sequence. Modern implementations further extend their capabilities to encode more complex geometric information, such as $2$D spatial (Dosovitskiy et al., 2021) or graph (Kreuzer et al., 2021) structures. For the analysis of any-variate time series, we distinguish between the temporal and variate structure. The temporal structure is equivalent to a sequential $1$D structure, such that we use standard $1$D sinusoidal embeddings $e^{\mathcal{T}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{T}}_{t}\in\mathbb{R}^{1\times D}$.
The variate structure exhibits great heterogeneity across domains. In domains with uni-variate and two-variate data, such as mono and stereo audio, the structure is either trivial or only requires a basic distinction between variates. In other domains, however, the variate structure may represent more complex relationships, such as $3$D manifolds for electroencephalography (EEG) or electrocardiography (ECG) data, or be of non-spatial nature, such as for financial data. Hence, we introduce learnable domain-specific variate embeddings to adequately address the heterogeneity across domains. These embeddings, denoted as $e^{\mathcal{V}}_{\mathcal{S}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$ for each variate $v$ in domain $S$, are designed to model the unique properties of a domain. They capture the inter-variate relationships and temporal dynamics specific to domain $S$, forming what can be considered as the signature of the very domain.
Finally, the patch, temporal, and domain-specific variate embeddings are summed to form the input token ${\bm{e}}_{v,t}=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}+\displaystyle{\bm{e}}^{\mathcal{T}}_{t}+\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$. These input tokens collectively constitute the final input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. To support batches of any-variate time series from multiple domains, we pad the variate dimension to the maximum number of variates in a batch $\overline{V}=\max_{S}V_{S}$. For samples where $V_{S}<\overline{V}$ or $T<\overline{T}$, attention masking is used to ensure that padded variate or temporal tokens are ignored. The domain-specific tokeniser is trained end-to-end with the Transformer layers.
Definition of (Sub-)Domains.
The domain-specific tokeniser is designed to integrate different datasets within a domain. Consider two EEG datasets, TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), which share $19$ identical variates but have different sampling frequencies of 500500500\,500Hz and 200200200\,200Hz, respectively. In this case, a single EEG-specific tokeniser ($V_{\text{EEG}}=19$) is sufficient to accommodate both sampling frequencies, i.e. ${\bm{E}}^{\mathcal{V}}_{\text{EEG-TDBrain}}=\displaystyle{\bm{E}}^{\mathcal{V}}_{\text{EEG-SEED}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},19}]^{\top}\in\mathbb{R}^{19\times D}$, as demonstrated in our experiments in Section 4. Note that while these positional embeddings are agnostic to variate ordering, we simplify processing by aligning the variate order across datasets within the same domain. Consider another EEG dataset, LEMON (Babayan et al., 2019), which includes $62$ electrodes. Of these, $15$ overlap with the electrodes in TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), while the remaining $47$ are unique to LEMON (Babayan et al., 2019). In this scenario, the EEG-specific tokeniser can be extended by the $47$ new variates ($V_{\text{EEG}}=66$), such that ${\bm{E}}^{\mathcal{V}}_{\text{EEG-LEMON}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},15},\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},20},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},66}]^{\top}\in\mathbb{R}^{62\times D}$. In this way, different datasets can be combined to approximate the underlying data distribution of a domain $S$, e.g. EEG, enabling the creation of large and diverse time series corpora.
Multi-Variate or Uni-Variate Analysis?
Consider the Electricity dataset (UCI, 2024), which contains electricity consumption data for $321$ households recorded from $2012$ to $2014$. These $321$ observations are sampled from an underlying population and are assumed to be independent and identically distributed (i.i.d.). In this scenario, we perform a uni-variate analysis ($V_{\text{Electricity}}=1$) of the data, initialising a single Electricity-specific variate embedding that models the hourly consumption of a household. In contrast, the Weather dataset (Wetterstation, 2024) contains $21$ climatological indicators, such as air temperature, precipitation, and wind speed, which are not i.i.d. because they directly interact and correlate with one another. Therefore, a multi-variate analysis ($V_{\text{Weather}}=21$) is conducted to account for the dependencies and interactions between the observations.
3.2 Pre-Training on Multi-Domain Time Series
We pre-train our model using masked data modelling (MDM) (He et al., 2022) to learn generalisable time series features across domains. We mask a subset of input tokens and only encode the visible (i.e. non-masked) tokens using an encoder $f(\cdot)$. Afterwards, we complement the encoded tokens with learnable mask tokens and feed them to a decoder $g(\cdot)$, reconstructing the original input tokens.
More precisely, we draw a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$, following the dual masking strategy proposed in Section 3.4, and apply it to the input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. Thus, we obtain a visible view ${\bm{E}}[\displaystyle{\bm{m}}]\in\mathbb{R}^{N_{1}\times D}$, where $N_{1}=\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}m_{v,t}$ and $N_{0}=(V_{S}\cdot T^{\prime})-N_{1}$ denote the number of visible and masked tokens, respectively. The visible view ${\bm{E}}[\displaystyle{\bm{m}}]$ is then fed to the encoder $f(\cdot)$ to compute the token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$:
${\bm{H}}=f(\displaystyle{\bm{E}}[\displaystyle{\bm{m}}]).$ (1)
To reconstruct the original input, these token features are fed to the decoder $g(\cdot)$ together with a special, learnable mask token ${\bm{e}}^{\mathcal{M}}\in\mathbb{R}^{1\times D}$, that is inserted at the masked positions where $m_{v,t}=0$:
${\bm{h}}^{\prime}_{v,t}=\begin{cases}\displaystyle{\bm{h}}_{v,t}&\text{if }m_{v,t}=1\\ \displaystyle{\bm{e}}^{\mathcal{M}}&\text{if }m_{v,t}=0\end{cases},$ (2)
such that ${\bm{H}}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. The decoder $g(\cdot)$ then predicts the reconstructed input $\widehat{{\bm{X}}}\in\mathbb{R}^{V_{S}\times(T^{\prime}\cdot P)}$:
$\widehat{{\bm{X}}}=g(\displaystyle{\bm{H}}^{\prime}),$ (3)
where $(T^{\prime}\cdot P)=\overline{T}$, i.e. the context length specified in time points. Eventually, the domain-specific tokeniser described in Section 3.1, the encoder $f(\cdot)$, and the decoder $g(\cdot)$ are optimised end-to-end using the mean squared error (MSE) loss on all reconstructed input tokens:
$\mathcal{L}_{\text{MSE}}=\frac{1}{V_{S}\cdot T^{\prime}}\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}\|\displaystyle{\bm{x}}_{v,t}-\displaystyle\widehat{{\bm{x}}}_{v,t}\|_{2}^{2}.$ (4)
3.3 Normalised Cross-Correlation Loss
MDM focuses on reconstructing masked parts of the data, emphasising local patterns through the MSE loss (4). However, time series often exhibit long-range dependencies, where past values influence future outcomes over extended periods. To accurately capture these global patterns, we introduce normalised cross-correlation (NCC) as a loss term in MDM for time series:
$\mathcal{L}_{\text{NCC}}=\frac{1}{V_{S}\cdot\overline{T}}\sum^{V_{S}}_{v=1}\sum^{\overline{T}}_{t=1}\hskip 2.84526pt\frac{1}{\sigma_{\displaystyle{\bm{x}}_{v}}\sigma_{\displaystyle\widehat{{\bm{x}}}_{v}}}(x_{v,t}-\mu_{\displaystyle{\bm{x}}_{v}})(\widehat{x}_{v,t}-\mu_{\displaystyle\widehat{{\bm{x}}}_{v}})\in\left[-1,1\right],$ (5)
where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively. Hence, to capture both local and global temporal dynamics, the total loss used to optimise OTiS is defined as
$\mathcal{L}=\mathcal{L}_{\text{MSE}}+\lambda\cdot\left(1-\mathcal{L}_{\text{NCC}}\right),$ (6)
where $\lambda$ is empirically set to $0.1$ during pre-training.
3.4 Dual Masking Strategy
We design the masking strategy to enhance foundational model capabilites in time series analysis. Specifically, we randomly select between two masking schemes during pre-training, namely random masking and post-fix masking. In 757575\,75of cases, we apply random masking, where each $m_{v,t}$ is independently sampled from a Bernoulli distribution with probability $p=1-\rho$, with $\rho$ denoting the masking ratio (i.e. $m_{v,t}\sim\text{Bernoulli}(1-\rho)$). This encourages the model to learn complex inter-variate relationships across the entire time series. In the remaining $25\,\%$ of cases, we employ post-fix masking, which masks the second half of the temporal dimension, leaving only the first half visible (i.e. $m_{v,t}=\mathbbm{1}_{[t\leq T^{\prime}/2]}$). The prediction of future values solely based on past observations simulates real-world forecasting conditions, helping the model to capture temporal causality. Overall, this dual masking strategy enables OTiS to learn both bidirectional relationships and temporal causality, which are essential for general time series analysis.
3.5 Fine-Tuning & Inference on (Unseen) Target Domains
Inclusion of Unseen Domains.
For a new domain $S$, a randomly initialised variate embedding ${\bm{E}}^{\mathcal{V}}_{\mathcal{S}}\in\mathbb{R}^{V_{S}\times D}$ is introduced. The domain-specific tokeniser is then fine-tuned alongside the encoder $f(\cdot)$, and, if required, the decoder $g(\cdot)$, for the specific downstream task, as described in the following.
Classification & Regression.
We use the encoder $f(\cdot)$ and the unmasked input sequence ${\bm{E}}$ to compute all token features ${\bm{H}}=f(\displaystyle{\bm{E}})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. We average-pool these features into a global token ${\bm{h}}^{*}\in\mathbb{R}^{1\times D}$, which we feed through a linear layer to obtain the final model prediction. We optimise a cross-entropy and MSE loss for the classification and regression tasks, respectively.
Forecasting.
We apply post-fix masking to generate a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$ for the forecasting task. The encoder $f(\cdot)$ is used to compute the visible token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$. We then concatenate the sequence with learnable mask tokens to form ${\bm{H}}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$, which is passed through the decoder $g(\cdot)$ to produce the final output. We optimise the MSE loss together with the NCC loss term over all reconstructed input tokens.
4 Experiments & Results
4.1 Model Variants and Implementation Details
We introduce OTiS in three different configurations, Base, Large, and Huge, with their specific architectures described in Appendix C.1, to explore scaling laws with respect to the model size. We set the patch size and stride to $P=24$, respectively, to split the time series into $T^{\prime}=\frac{\overline{T}}{P}$ non-overlapping patches along the time dimension. For pre-training, the context length specified in time points is set to $\overline{T}=1008$, resulting in $T^{\prime}=42$ sinusoidal temporal embeddings. If longer context lengths are required during fine-tuning, these embeddings are linearly interpolated (i.e. $T^{\prime}_{\text{ft}}\geq 42$) to offer greater flexibility for downstream applications. We tune the hyperparameters for pre-training and fine-tuning as described in Appendix C. An overview of the computational costs is provided in Appendix D.
4.2 Large and Diverse Pre-Training Corpus
We aim to develop a general time series model that fully handles the heterogeneity in real-world data. Specifically, our model is designed to handle time series with different variate counts $V_{S}$, inter-variate relationships, temporal dynamics, and sampling frequency, ensuring flexibility for downstream tasks. To this end, we pre-train our model on a large and diverse corpus of publicly available data spanning $8$ domains, with a total of $640,187$ samples and $11$ billion time points, as summarised in Table 1. A detailed description of the datasets included in our pre-training corpus can be found in Appendix A. The time series corpus is split into $612,394$ training and $27,793$ validation samples for pre-training.
Table 1: Overview of our large and diverse pre-training corpus. The corpus is built with unlabelled data from eight domains, encompassing medicine, engineering, natural sciences, and finance.
Domain Name Samples Variates Time points Frequency Disk size
$S$ $V_{S}$
ECG MIMIC-IV-ECG 2023 $400,000$ $12$ $5,000$ 500500500\,500Hz 909090\,90GB
Temperature DWD 2024 $203,340$ $1$ $720$ (hourly) 278278278\,278mHz 614614614\,614MB
Audio (stereo) AudioSet-$20$K 2017 $16,123$ $2$ $441,000$ 44.144.144.1\,44.1kHz 535353\,53GB
Audio (mono) AudioSet-$20$K 2017 $3,491$ $1$ $441,000$ 44.144.144.1\,44.1kHz 666\,6GB
Electromechanics FD-A 2016 $13,640$ $1$ $5,120$ 646464\,64kHz 161161161\,161MB
EEG TDBrain 2022 $2,692$ $19$ $60,000$ 500500500\,500Hz 121212\,12GB
EEG SEED 2015 $675$ $19$ $37,000$ 200200200\,200Hz 222\,2GB
Banking NN5555 2012 $111$ $1$ $971$ (daily) $12\,\mu$Hz 370370370\,370KB
Economics FRED-MD 2016 $107$ $1$ $728$ (monthly) 386386386\,386nHz 330330330\,330KB
Economics Exchange 2018 $8$ $1$ $7,588$ (daily) $12\,\mu$Hz 240240240\,240KB
640,187 11,052,756,981 164 GB
4.3 Benchmarking Across Domains and Tasks
To assess the utility of OTiS in real-world settings, we conduct experiments on three key use cases in time series analysis: classification, regression, and forecasting. For classification, we perform binary epilepsy detection using EEG (Epilepsy 2001), multi-class fault detection in rolling bearings from vibration signals (FD-B 2016), multi-class hand-gesture classification with accelerometer signals (Gesture 2009), and multi-class muscular disease classification using electromyographie (EMG 2000). In regression, we predict five imaging-derived cardiac phenotypes from $12$-lead ECG (LVEDV, LVESV, LVSV, LVEF, LVM 2020). For forecasting, we predict electricity transformer temperature (ETT 2021), weather (Weather 2024), and electricity consumption (Electricity 2024). Further dataset details are provided in Appendix B. We adhere to the established data splitting and evaluation procedures for the classification (Zhang et al., 2022), regression (Turgut et al., 2023), and forecasting (Zhou et al., 2021) tasks. All experiments are reported as mean and standard deviation across five seeds set during fine-tuning.
Table 2: Classification and regression performance on a total of $9$ benchmark tasks. OTiS is competitive with specialised baselines, setting new state-of-the-art on $6$ tasks and even outperforming the multimodal CM-AE and MMCL. This demonstrates the capability of OTiS to extract high-level semantics. Best score in bold, second best underlined. indicates tasks in previously unseen domains.
(a) Classification [Accuracy (ACC1$\uparrow$) in %]
Model Epilepsy FD-B Gesture EMG
SimCLR 2020 90.71 49.17 48.04 61.46
CoST 2022 88.40 47.06 68.33 53.65
TS2Vec 2022 93.95 47.90 69.17 78.54
TF-C 2022 94.95 69.38 76.42 81.71
Ti-MAE 2023 89.71 60.88 71.88 69.99
SimMTM 2024 95.49 69.40 80.00 97.56
OTiS-Base 94.25 99.24 63.61 97.56
OTiS-Large 94.03 98.62 62.50 98.37
OTiS-Huge 91.48 98.32 63.61 98.37
(b) Regression [R-squared ($R^{2}$uparrow)]
Model LVEDV LVESV LVSV LVEF LVM
ViT 2023 0.409 0.396 0.299 0.175 0.469
MAE 2023 0.486 0.482 0.359 0.237 0.573
CM-AE* 2023 0.451 0.380 0.316 0.103 0.536
MMCL* 2023 0.504 0.503 0.370 0.250 0.608
OTiS-Base 0.509 0.512 0.391 0.292 0.592
OTiS-Large 0.504 0.503 0.371 0.267 0.592
OTiS-Huge 0.505 0.510 0.376 0.281 0.593
* Models that incorporate paired imaging data during pre-training.
OTiS demonstrates strong classification capabilities, setting two new benchmarks, as shown in Table 2(a). Additionally, OTiS excels in predicting imaging-derived cardiac phenotypes, even surpassing multimodal baselines that incorporate imaging data during pre-training, as summarised in Table 2(b). Our model outperforms the baselines in $4$ out of $5$ regression tasks. As shown in Table 3, OTiS also exhibits strong forecasting capabilities, outperforming current models in $4$ out of $6$ benchmarks. Notably, all of the forecasting tasks are performed in previously unseen domains. A visualisation of the forecast predictions can be found in Appendix F. Overall, OTiS outperforms both specialised and general state-of-the-art baselines in $10$ out of $15$ diverse applications across $8$ domains, demonstrating its strong utility and generalisability across (unseen) domains and tasks.
Table 3: Forecasting performance on $6$ benchmark tasks. OTiS is competitive with specialised and general baselines, setting new state-of-the-art on $4$ tasks and showcasing its ability to capture local time series features. A forecasting horizon of $96$ time points is predicted from the past $336$ (*$512$, ${}^{+}904$) time points. Mean squared error (MSE $\downarrow$) is reported. Best score in bold, second best underlined. indicates tasks in previously unseen domains.
Model ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
N-BEATS 2019 0.399 0.327 0.318 0.197 0.152 0.131
Autoformer 2021 0.435 0.332 0.510 0.205 0.249 0.196
TimesNet 2022a 0.384 0.340 0.338 0.187 0.172 0.168
DLinear 2023 0.375 0.289 0.299 0.167 0.176 0.140
PatchTST 2023 0.370 0.274 0.293 0.166 0.149 0.129
Time-LLM 2023 0.408 0.286 0.384 0.181
GPT4TS 2023 0.376 0.285 0.292 0.173 0.162 0.139
MOMENT* 2024 0.387 0.288 0.293 0.170 0.154 0.136
MOIRAI+ 2024 0.375 0.277 0.335 0.189 0.167 0.152
OTiS-Base 0.424 0.212 0.337 0.161 0.139 0.128
OTiS-Large 0.446 0.205 0.362 0.173 0.142 0.127
OTiS-Huge 0.461 0.215 0.384 0.181 0.149 0.132
Experiments could not be conducted on a single NVIDIA RTX A6000600060006000-$48$GB GPU.
4.4 Domain Signature Analysis
A key component of OTiS is its use of domain-specific variate embeddings. While these embeddings are randomly initialised, we expect them to capture unique domain characteristics during training, eventually serving as the signature of their respective domain. To validate this hypothesis, we analyse the domain-specific variate embeddings after pre-training using principal component analysis (PCA).
First, we find that OTiS unifies time series from diverse domains into a meaningful latent space, where embeddings of domains with shared high-level semantics cluster together, as depicted in Appendix E.1. For example, embeddings of mono and stereo audio group closely, as do those of banking and economics. Moreover, EEG-specific embeddings are clearly separated and ECG-specific embeddings form a tight cluster.
Second, OTiS preserves low-level semantics specific to each domain, such as the relationships between variates. To explore this, we focus on the learned variate embeddings of EEG, the most complex domain in our corpus. EEG variates correspond to actual electrodes, each associated with a $3$D position in space or a $2$D position on the scalp, which is ideal for studying inter-variate relationships. Our analysis covers both (i) variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during multi-domain pre-training, and (ii) variate embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning of the pre-trained OTiS. We find that the first three principal components explain (i) $74.7\,\%$ and (ii) $87.9\,\%$ of the variance, indicating that the $3$D PCA projections of the variate embeddings correspond closely to the actual electrode positions. A visualisation of the learned $3$D electrode layout is provided in Appendix E. Additionally, we linearly align these positions with the true EEG electrode layout on the $2$D scalp manifold, as depicted in Figure 3. The alignment yields $R^{2}$ values of (i) $0.81$ and (ii) $0.95$, confirming a strong correspondence between the learned variate embeddings and the actual electrode layout.
Figure 3: $2$D PCA projections of the learned EEG-specific variate embeddings, overlaid on the actual EEG electrode layout. (Left) Embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. (Right) Embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning. The embeddings accurately reflect the spatial electrode layout, as confirmed by high correlations (R2) between the PCA projections $\bullet$ and the true layout $\circ$.
4.5 Scaling Study
Figure 4: Performance of OTiS with different numbers of pre-training samples. Shaded regions indicate the standard deviation across $5$ seeds. Increasing dataset size generally improves downstream performance. Scaling model size requires even larger pre-training corpora to be effective.
We analyse the scaling behaviour of OTiS with respect to model and dataset size. To this end, we subsample the pre-training data to $10\,\%$ and $1\,\%$ of its original size, ensuring that each subset is fully contained within the corresponding superset. We evaluate the downstream performance of all OTiS variants across classification, regression, and forecasting tasks, as depicted in Figure 4.
The experiments demonstrate that downstream performance generally scales with dataset size, achieving the best results with the full pre-training dataset. This trend, however, does not directly apply to model size, which is in line with the scaling behaviour observed in current time series foundational models (Woo et al., 2024; Goswami et al., 2024). Given that performance generally improves across all models with increasing data size, we hypothesise that scaling the model size could prove beneficial with even larger pre-training corpora.
4.6 Ablation Study
Figure 5: Ablation study on key components of OTiS. Downstream performance is analysed across $5$ seeds. A leave-one-out approach is used to evaluate the influence of each component. The default setting, which includes all components, demonstrates superior model capabilities across tasks.
We perform an ablation study to analyse the impact of OTiS’ key components: the domain-specific tokeniser, dual masking strategy, and normalised cross-correlation (NCC) loss. As shown in Figure 5, the best and most robust performance is achieved when all components are used during pre-training.
Replacing the domain-specific variate embeddings with domain-agnostic embeddings (i.e. learnable embeddings shared across all domains) consistently led to inferior performance across all tasks, demonstrating the importance of capturing domain-specific data characteristics during tokenisation. Switching from dual masking to random masking resulted in performance degradation, although the impact was less notable for generative tasks than for discriminative tasks. We hypothesise that the NCC loss already captures temporal causality, which is particularly crucial for generative tasks like forecasting. Overall, removing the NCC loss caused performance declines across all downstream tasks, emphasising the role of long-range dependencies for general time series understanding.
5 Discussion & Conclusion
In this study, we explore the problem of effective pre-training on heterogeneous time series corpora. Time series vary substantially across domains, e.g. with respect to inter-variate relationships and temporal dynamics, rendering generalisable feature extraction from multi-domain time series difficult. To address this issue, we present OTiS, an open model for general time series analysis, specifically designed to handle multi-domain heterogeneity. Our novel multi-domain pre-training paradigm, including a domain-specific tokeniser with learnable signatures, a dual masking strategy, and a normalised cross-correlation (NCC) loss, enables OTiS to extract generalisable time series features.
In extensive experiments, we demonstrate that OTiS generalises well across $15$ diverse downstream applications spanning $8$ distinct domains, achieving competitive performance with both specialised and general state-of-the-art (SOTA) models. In a qualitative analysis, we further show that OTiS unifies time series from diverse domains in a meaningful latent space, while preserving low-level semantics of a domain including the inter-variate relationships. Thereby, our work establishes a strong foundation for future advancements in interpretable and general time series analysis.
Limitations.
While OTiS outperforms SOTA models across $10$ tasks, our experiments in low-data regimes suggest that larger pre-training corpora could further enhance its performance. Unlike in NLP and CV, where large datasets are curated from web-crawled data, foundational models in time series, including OTiS, still rely on manually curated datasets. Future work could explore fully automatic pipelines, e.g. using embedding similarity, to filter and rebalance multi-domain time series from the web. OTiS could further benefit from processing domain signatures during inference, potentially unlocking zero-shot capabilities, similarly to those seen in foundational models in NLP and CV.
Appendix A Large Multi-Domain Pre-Training Corpus
In this section, we present an overview of our large and diverse pre-training corpus. The corpus consists of publicly available data spanning eight domains, with a total of $640,187$ samples and $11$ billion time points. In the following, we provide a detailed breakdown of the domains and the datasets they encompass. Note that we apply channel-wise standard normalisation to the datasets unless otherwise specified.
ECG.
The MIMIC-IV-ECG dataset (Gow et al., 2023) contains diagnostic $10$-second, $12$-lead ECG recordings sampled at a frequency of $500$ Hz. While the entire dataset comprises $800,035$ samples, we include only the first half of the recordings available in the database, preventing the ECG data from predominating in the pre-training corpus. To remove the baseline drift from the ECG data, we use the asymmetric least square smoothing technique (Zhang et al., 2010). Note that we apply standard normalisation separately to the Einthoven, Goldberger, and Wilson leads.
Temperature.
The Deutscher Wetterdienst (DWD) dataset (Wetterdienst, 2024) contains hourly air temperature measurements from $629$ weather stations across Germany. Since the recording length varies significantly, ranging from $763$ to $1,148,290$ hours per station, we split the data into chunks of $720$ hours (approximately one month).
Audio.
The AudioSet dataset (Gemmeke et al., 2017) contains $10$-second YouTube clips for audio classification, featuring $527$ types of audio events that are weakly annotated for each clip. The full training set includes a class-wise balanced subset (AudioSet-$20$K, $22,176$ clips) and an unbalanced (AudioSet-$2$M $2,042,985$ clips) set. For our pre-training corpus, we use the balanced AudioSet-$20$K, which contains $3,491$ mono and $16,123$ stereo recordings, all sampled at $44,100\,$Hz.
Electromechanics.
The FD-A dataset (Lessmeier et al., 2016) collects vibration signals from rolling bearings in a mechanical system for fault detection purposes. Each sample consists of $5,120$ timestamps, indicating one of three mechanical device states. Note that the FD-B dataset is similar to FD-A but includes rolling bearings tested under different working conditions, such as varying rotational speeds.
EEG.
The TDBrain dataset (Van Dijk et al., 2022) includes raw resting-state EEG data from $1,274$ psychiatric patients aged $5$ to $89$, collected between $2001$ and $2021$. The dataset covers a range of conditions, including Major Depressive Disorder ($426$ patients), Attention Deficit Hyperactivity Disorder ($271$ patients), Subjective Memory Complaints ($119$ patients), and Obsessive-Compulsive Disorder ($75$ patients). The data was recorded at 500500500\,500Hz using $26$ channel EEG-recordings, based on the $10$-$10$ electrode international system.
The SEED dataset (Zheng & Lu, 2015) contains EEG data recorded under three emotional states: positive, neutral, and negative. It comprises EEG data from $15$ subjects, with each subject participating in experiments twice, several days apart. The data is sampled at 200200200\,200Hz and recorded using $62$ channel EEG-recordings, based on the $10$-$20$ electrode international system.
For simplicity, we only consider the $19$ channels common to both datasets, i.e. the channels that correspond to the $10$-$20$ electrode international system.
Banking.
The NN5555 competition dataset (Taieb et al., 2012) consists of daily cash withdrawals observed at $111$ randomly selected automated teller machines across various locations in England.
Economics.
The FRED-MD dataset (McCracken & Ng, 2016) contains $107$ monthly time series showing a set of macro-economic indicators from the Federal Reserve Bank of St Louis. The data was extracted from the FRED-MD database.
The Exchange dataset (Lai et al., 2018) records the daily exchange rates of eight different nations, including Australia, Great Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, ranging from $1990$ to $2016$.
Appendix B Benchmark Dataset Details
We provide an overview of the datasets used to benchmark our model on $15$ distinct downstream applications in Table 4. Our benchmarking experiments include classifcation, regression, and forecasting tasks.
Table 4: Experimental setup with all conducted downstream applications, evaluation metrics, and datasets used for benchmarking.
Task Metric Dataset
Domain Name Samples Variates Time points Frequency
$S$ $V_{S}$
Classification ACC EEG Epilepsy 2001 $11,500$ $1$ $178$ 174174174\,174Hz
Electromechanics FD-B 2016 $13,640$ $1$ $5,120$ 646464\,64kHz
Acceleration Gesture 2009 $560$ $3$ $206$ 100100100\,100Hz
EMG EMG 2000 $204$ $1$ $1,500$ 444\,4kHz
Regression $R^{2}$ ECG UK BioBank 2015 $18,926$ $12$ $5,000$ 500500500\,500Hz
Forecasting MSE Energy ETTh1 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
ETTh2 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
ETTm1 2021 $1$ $7$ $69,680$ (minutely) 1.11.11.1\,1.1mHz
ETTm2 2021 $1$ $7$ $69,680$ (minutely) 1.11.11.1\,1.1mHz
Weather Weather 2024 $1$ $21$ $52,696$ (minutely) 2.82.82.8\,2.8mHz
Electricity Electricity 2024 $321$ $1$ $26,304$ (hourly) 278278278\,278mHz
Appendix C Experiment Details
C.1 Model Variants
To explore the scaling laws with respect to the model size, we provide OTiS in three variants, as summerised in Table 5.
Table 5: Details of model variants.
Model Layers Hidden size D MLP size Heads $d_{kv}$ Parameters
OTiS-Base $12$ $192$ $768$ $3$ $64$ 888\,8M
OTiS-Large $18$ $384$ $1536$ $6$ $64$ 444444\,44M
OTiS-Huge $24$ $576$ $2304$ $8$ $72$ 131131131\,131M
C.2 Pre-Training & Fine-Tuning Parameters
We provide the hyperparameters used to pre-train all variants of OTiS in Table 6. The hyperparameters used to fine-tune our models for the classification, regression, and forecasting tasks are provided in Table 7, 8, and 9, respectively.
Table 6: Hyperparameters used for pre-training. Pre-training is performed on $4$ NVIDIA A100100100100-$80$GB GPUs. A cosine learning rate scheduler is applied with a $10\,\%$ warmup. All OTiS configurations use a shallow decoder with 222\,2M parameters, consisting of $4$ layers with a hidden size of $160$, an MLP with size $640$, and $5$ heads.
Model Epochs Batch size Base LR LR decay NCC $\lambda$ Mask ratio $\rho$ Weight decay
OTiS-Base $200$ $5120$ $3$e-$5$ cosine $0.1$ $0.75$ $0.10$
OTiS-Large $200$ $3328$ $1$e-$5$ cosine $0.1$ $0.75$ $0.15$
OTiS-Huge $200$ $2880$ $3$e-$6$ cosine $0.1$ $0.75$ $0.05$
Table 7: Hyperparameters used for fine-tuning the classification tasks on a single NVIDIA RTX A6000600060006000-$48$GB GPU. A cosine learning rate scheduler is applied with a $10\,\%$ warmup.
Dataset Model Epochs Batch size Base LR Drop path Layer decay Weight decay Label smoothing
Epilepsy OTiS-Base $75$ $32$ $1$e-$3$ $0.2$ $0.75$ $0.2$ $0.1$
OTiS-Large $75$ $32$ $3$e-$3$ $0.2$ $0.50$ $0.1$ $0.1$
OTiS-Huge $75$ $32$ $3$e-$3$ $0.0$ $0.75$ $0.2$ $0.2$
FD-B OTiS-Base $75$ $32$ $3$e-$4$ $0.0$ $0.75$ $0.1$ $0.1$
OTiS-Large $75$ $32$ $1$e-$3$ $0.1$ $0.75$ $0.1$ $0.2$
OTiS-Huge $75$ $32$ $3$e-$4$ $0.1$ $0.75$ $0.2$ $0.1$
Gesture OTiS-Base $75$ $32$ $1$e-$2$ $0.1$ $0.25$ $0.2$ $0.1$
OTiS-Large $75$ $32$ $3$e-$3$ $0.2$ $0.75$ $0.1$ $0.0$
OTiS-Huge $75$ $32$ $1$e-$2$ $0.0$ $0.75$ $0.1$ $0.1$
EMG OTiS-Base $75$ $32$ $1$e-$3$ $0.2$ $0.75$ $0.1$ $0.2$
OTiS-Large $75$ $32$ $3$e-$3$ $0.1$ $0.75$ $0.2$ $0.1$
OTiS-Huge $75$ $32$ $3$e-$3$ $0.1$ $0.75$ $0.2$ $0.2$
Table 8: Hyperparameters used for fine-tuning the regression tasks on a single NVIDIA RTX A6000600060006000-$48$GB GPU. A cosine learning rate scheduler is applied with a $10\,\%$ warmup.
Dataset Model Epochs Batch size Base LR Drop path Layer decay Weight decay
UK BioBank OTiS-Base $50$ $192$ $3$e-$4$ $0.2$ $0.75$ $0.1$
OTiS-Large $50$ $160$ $1$e-$4$ $0.2$ $0.75$ $0.1$
OTiS-Huge $50$ $200$ $1$e-$4$ $0.2$ $0.75$ $0.1$
Table 9: Hyperparameters used for fine-tuning the forecasting tasks. A cosine learning rate scheduler is applied with a $10\,\%$ warmup.
Dataset Model Epochs Batch size Base LR NCC $\lambda$ Weight decay
ETTh1 OTiS-Base $1000$ $1$ $1$e-00 $0.1$ $0.15$
OTiS-Large $1000$ $1$ $1$e-$1$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $3$e-$1$ $0.1$ $0.15$
ETTh2 OTiS-Base $1000$ $1$ $1$e-00 $0.2$ $0.25$
OTiS-Large $1000$ $1$ $1$e-$1$ $0.1$ $0.25$
OTiS-Huge $1000$ $1$ $3$e-$1$ $0.0$ $0.25$
ETTm1 OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.1$ $0.15$
ETTm2 OTiS-Base $1000$ $1$ $3$e-$1$ $0.1$ $0.25$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.25$
Weather OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.05$
Electricity OTiS-Base $250$ $32$ $3$e-$2$ $0.0$ $0.25$
OTiS-Large $250$ $32$ $3$e-$2$ $0.0$ $0.15$
OTiS-Huge $250$ $32$ $3$e-$2$ $0.2$ $0.15$
Appendix D Computation Costs
We provide an overview of the computational resources used to train OTiS in Table 10.
Table 10: Computational resources used to pre-train OTiS. Note that fine-tuning and inference of all OTiS variants on downstream applications were performed using a single NVIDIA RTX A6000600060006000-$48$GB and $32$ CPUs.
Model Parameters Power consumption CPU count GPU
Count Hours Type
OTiS-Base $8$ M 700700700\,700W* $128$ $4$ $115$ NVIDIA A100100100100-$80$GB
OTiS-Large $44$ M 800800800\,800W* $128$ $4$ $154$ NVIDIA A100100100100-$80$GB
OTiS-Huge $131$ M 960960960\,960W* $128$ $4$ $219$ NVIDIA A100100100100-$80$GB
* Total power consumption across all GPUs.
Total hours across all GPUs.
Appendix E Domain Signature Analysis
To analyse the domain signatures, we reduce the dimensionality of the domain-specific variate embeddings by employing a principal component analysis (PCA). Our analysis shows that OTiS unifies time series from diverse domains into a meaningful latent space, while accurately capturing the inter-variate relationships within a domain.
E.1 Inter-Domain Analysis
A visualisation of all domain-specific variate embeddings learned during pre-training is provided in Figure 6. We find that OTiS learns a meaningful latent space, where embeddings of domains with shared high-level semantics cluster closely together.
(a) $3$D projection.
(b) $2$D projection.
Figure 6: PCA projections of the domain-specific variate embeddings learned during pre-training. OTiS unifies time series from diverse domains in a meaningful latent space, while correctly encoding the inter-variate relationships within a domain. Mono ($\bullet$) and stereo ($\bullet$) audio-specific embeddings cluster closely together, as do those for banking ($\bullet$) and economics ($\bullet$). Clear separation is observed for EEG-specific embeddings ($\bullet$), while also ECG-specific embeddings ($\bullet$) form a tight cluster.
E.2 Intra-Domain Analysis
A visualisation of the domain-specific variate embeddings for EEG and ECG are given in Figure 7 and Figure 8, respectively. We find that OTiS preserves low-level semantics specific to each domain, accurately capturing the relationships between variates.
(a) True electrode layout.
(b) Implicitly learned electrode layout.
Figure 7: $3$D PCA projections of the variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. The labels of each node correspond to the electrode names.
(a) Side view.
(b) Bird’s-eye view.
(c) Top view.
Figure 8: $3$D PCA projections of the variate embeddings for standard $12$-lead ECG learned during pre-training. The embeddings accurately reflect the true physiological structure of the ECG leads. The V1111-V6666 leads, arranged on the rib cage from the sternum to the mid-axillary line, represent a $3$D view of the human heart. In contrast, the I-II-III leads and aVR-aVL-aVF leads, derived from electrodes placed on one foot and both arms, form a planar $2$D triangle.
Appendix F Forecast Visualisation
We visualise the performance of our model on $6$ forecasting benchmarks in Figure 9.
Figure 9: Visualisation of OTiS-Base forecast predictions on $6$ benchmark datasets. A forecasting horizon of $96$ time points is predicted from the past $336$ time points. Ground truth in blue, prediction in orange. Areas highlighted in grey are not visible to the model.

### v2

@@
 Abstract
-In natural language processing and computer vision, self-supervised pre-training on large datasets unlocks foundational model capabilities across domains and tasks.
-However, this potential has not yet been realised in time series analysis, where existing methods disregard the heterogeneous nature of time series characteristics.
-Time series are prevalent in many domains, including medicine, engineering, natural sciences, and finance, but their characteristics vary significantly in terms of variate count, inter-variate relationships, temporal dynamics, and sampling frequency.
-This inherent heterogeneity across domains prevents effective pre-training on large time series corpora.
-To address this issue, we introduce OTiS, an open model for general time series analysis, that has been specifically designed to handle multi-domain heterogeneity.
-We propose a novel pre-training paradigm including a tokeniser with learnable domain-specific signatures, a dual masking strategy to capture temporal causality, and a normalised cross-correlation loss to model long-range dependencies.
-Our model is pre-trained on a large corpus of $640,187$ samples and $11$ billion time points spanning $8$ distinct domains, enabling it to analyse time series from any (unseen) domain.
-In comprehensive experiments across $15$ diverse applications - including classification, regression, and forecasting - OTiS showcases its ability to accurately capture domain-specific data characteristics and demonstrates its competitiveness against state-of-the-art baselines.
-Our code and pre-trained weights are publicly available at https://github.com/oetu/otis.
+Recent breakthroughs in natural language processing and computer vision, driven by efficient pre-training on large datasets, have enabled foundation models to excel on a wide range of tasks.
+However, this potential has not yet been fully realised in time series analysis, as existing methods fail to address the heterogeneity in large time series corpora.
+Prevalent in domains ranging from medicine to finance, time series vary substantially in characteristics such as variate count, inter-variate relationships, temporal patterns, and sampling frequency.
+To address this, we introduce a novel pre-training paradigm specifically designed to handle time series heterogeneity.
+We propose a tokeniser with learnable domain signatures, a dual masking strategy, and a normalised cross-correlation loss, enabling our open model for general time series analysis (OTiS) to efficiently learn from large time series corpora.
+Extensive benchmarking on diverse tasks, such as classification, regression, and forecasting, demonstrates that OTiS outperforms state-of-the-art baselines.
+Our code and pre-trained weights are available at https://github.com/oetu/otis.
+Machine Learning, ICML
 1 Introduction
+Figure 1: Open model for general time series analysis (OTiS).
+Its tokeniser accounts for varying time series characteristics across domains, such as distinct variate counts, inter-variate relationships, temporal patterns, and sampling frequencies.
+It can be fine-tuned on limited data from any domain, including previously unseen ones, to perform classification, regression, and forecasting tasks.
 In natural language processing (NLP) or computer vision (CV), generalisable language features, e.g.
@@
 edges and shapes (Geirhos et al., 2019; Dosovitskiy et al., 2021; Oquab et al., 2024), are learned from large-scale data.
-Self-supervised pre-training paradigms are designed to account for the specific properties of language (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023) or imaging (Zhou et al., 2022; Cherti et al., 2023; Oquab et al., 2024), unlocking foundational model capabilities that apply to a wide range of domains and downstream tasks.
-This potential, however, remains largely unrealised in time series due to the lack of self-supervised pre-training paradigms that account for the heterogeneity of time series across domains.
+Self-supervised pre-training paradigms are designed to account for the specific properties of language (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023) or imaging (Zhou et al., 2022; Cherti et al., 2023; Oquab et al., 2024), unlocking foundational model capabilities that apply to a wide range of tasks.
+This potential, however, remains largely unrealised in time series due to the lack of pre-training paradigms that effectively address time series heterogeneity across domains.
 Time series are widespread in everyday applications and play an important role in various domains, including medicine (Pirkis et al., 2021), engineering (Gasparin et al., 2022), natural sciences (Ravuri et al., 2021), and finance (Sezer et al., 2020).
-They differ substantially with respect to the number of variates, inter-variate relationships, temporal dynamics, and sampling frequency (Fawaz et al., 2018; Ismail Fawaz et al., 2019; Ye & Dai, 2021; Wickstrøm et al., 2022).
+They differ substantially with respect to the number of variates, inter-variate relationships, temporal patterns, and sampling frequency (Fawaz et al., 2018; Ismail Fawaz et al., 2019; Ye & Dai, 2021; Wickstrøm et al., 2022).
 For instance, standard $10$-$20$ system electroencephalography (EEG) recordings come with up to $256$ variates (Jurcak et al., 2007), while most audio recordings have only $1$ (mono) or $2$ (stereo) variates.
 Weather data shows high periodicity, whereas financial data is exposed to long-term trends.
-Both domains encompass low-frequency data recorded on an hourly (278278278\,278mHz), daily ($12\,\mu$Hz), or even monthly (386386386\,386nHz) basis, while audio data is sampled at high frequencies of 44.144.144.1\,44.1kHz or more.
+Both domains encompass low-frequency data recorded on an hourly ($278\,\mu$Hz), daily ($12\,\mu$Hz), or even monthly (386386386\,386nHz) basis, while audio data is sampled at high frequencies of 44.144.144.1\,44.1kHz or more.
 Overall, this heterogeneity across domains renders the extraction of generalisable time series features difficult (Fawaz et al., 2018; Gupta et al., 2020; Iwana & Uchida, 2021; Ye & Dai, 2021).
-While most existing self-supervised pre-training methods for time series are limited to single domains (Wu et al., 2021; 2022a; Nie et al., 2023; Dong et al., 2024; Jiang et al., 2024), recent works propose simple techniques to incorporate time series from multiple domains (Yang et al., 2024; Das et al., 2024; Woo et al., 2024; Liu et al., 2024).
-These works for instance crop all time series into segments of unified size (Jiang et al., 2024), resample them to a uniform frequency (Yang et al., 2024), or analyse each variate of a multi-variate time series independently (Liu et al., 2024).
-While these naive techniques address differences in sampling frequency and variate count, they degrade the original time series and neglect the critical inter-variate relationships and temporal dynamics required for effective real-world analysis.
-Consequently, there is a clear need for pre-training strategies that adequately handle heterogeneity in time series to unlock foundational model capabilities.
-In this work, we propose a novel multi-domain pre-training paradigm that addresses the full spectrum of time series heterogeneity across domains.
-Our approach facilitates the comprehensive extraction of generalisable features from diverse time series.
-Pre-trained on a large corpus of publicly available data, our open model for general time series analysis (OTiS) can be fine-tuned on limited data of any (unseen) domain to perform a variety of downstream tasks, as showcased in Figure 1.
-Figure 1: Overview of OTiS.
-Pre-trained on a large corpus of time series from diverse domains, OTiS enables general time series analysis.
-Its domain-specific tokeniser addresses time series heterogeneity across domains - including different numbers of variates, inter-variate relationships, temporal dynamics, and sampling frequencies - by learning unique domain signatures.
-After pre-training, the model can be fine-tuned on limited data from any domain, including previously unseen ones, to perform various tasks such as classification, regression, and forecasting.
+While most existing self-supervised pre-training methods for time series are limited to single domains (Wu et al., 2021, 2022a; Nie et al., 2023; Dong et al., 2024; Jiang et al., 2024), recent works propose simple techniques to incorporate time series from multiple domains (Yang et al., 2024; Das et al., 2024; Woo et al., 2024; Liu et al., 2024).
+These works, for instance, crop all time series into segments of unified size (Jiang et al., 2024), resample them to a uniform frequency (Yang et al., 2024), or analyse each variate of a multi-variate time series independently (Liu et al., 2024).
+While these naive techniques address domain differences in sampling frequency and variate count, they degrade the original time series and fail to account for inter-variate relationships and temporal patterns, which vary heavily across domains.
+Consequently, there is a clear need for pre-training strategies that effectively address data heterogeneity across domains.
+In this work, we propose a novel multi-domain pre-training paradigm that accounts for the varying characteristics across domains.
+This approach enables our open model for general time series analysis (OTiS) to be pre-trained on a large and diverse time series corpus, unlocking foundational capabilities.
 Our key contributions can be summarised as follows:
   1. 1.
-    We present OTiS, an open model for general time series analysis, with our entire pipeline and pre-trained weights publicly available at https://github.com/oetu/otis.
+    We propose a novel multi-domain pre-training paradigm that enables the learning of generalisable time series features.
+We introduce a novel tokeniser to account for the varying characteristics across domains, a dual masking strategy to capture bidirectional relationships and temporal causality, and a normalised cross-correlation loss to learn global temporal patterns.
   2. 2.
-    We propose a novel pre-training paradigm based on masked data modelling to address heterogeneity in multi-domain time series.
-Our approach includes a novel tokeniser with learnable signatures to capture domain-specific data characteristics, a dual masking strategy to learn temporal causality, and a normalised cross-correlation loss to model long-range dependencies.
+    We pre-train OTiS on a large corpus spanning $8$ domains from medicine, engineering, natural sciences, and finance.
+With $640,187$ samples and $11$ billion time points, it covers diverse time series characteristics, essential for developing foundational model capabilities.
   3. 3.
-    We pre-train OTiS on a large corpus of publicly available time series from $8$ domains, spanning medicine, engineering, natural sciences, and finance.
-With $640,187$ samples and $11$ billion time points, this corpus represents diverse time series characteristics, enabling generalisable feature extraction.
-  4. 4.
-    We evaluate OTiS across $15$ downstream applications, including classification, regression, and forecasting.
-Our comprehensive analysis demonstrates that OTiS accurately captures domain-specific data characteristics and is competitive with both specialised and general state-of-the-art (SOTA) models, achieving new SOTA performance in $10$ tasks.
-Notably, none of the baselines is capable of performing all the tasks covered by OTiS.
+    Evaluated across $15$ applications, OTiS is competitive with specialised and general state-of-the-art models, setting new benchmarks in $10$ tasks.
+Notably, none of the baselines can handle the variety of tasks covered by OTiS, i.e.
+classification, regression, and forecasting.
 2 Related Works
 2.1 Self-Supervised Learning for Time Series
-Time series vary significantly across domains, with differences in the number of variates, inter-variate relationships, temporal dynamics, and sampling frequencies.
+Time series vary significantly across domains, with differences in the number of variates, inter-variate relationships, temporal patterns, and sampling frequencies.
 Due to this inherent heterogeneity, most existing works focus on pre-training models within a single domain (Oreshkin et al., 2019; Tang et al., 2020; Wu et al., 2021; Zhou et al., 2021; Wu et al., 2022a; Woo et al., 2022; Yue et al., 2022; Zhang et al., 2022; Li et al., 2023; Nie et al., 2023; Zeng et al., 2023; Dong et al., 2024).
-To develop more generalisable time series models, recent methods have explored multi-domain pre-training by addressing certain aspects of the heterogeneity, such as differences in variate count and sampling frequency.
-For instance, Liu et al.
-(2024) treat each variate in multi-variate time series independently to standardise generative tasks like forecasting, while Goswami et al.
+To develop more general time series models, recent works have explored multi-domain pre-training by partially addressing this heterogeneity, including the different numbers of variates and sampling frequencies.
+Liu et al.
+(2024) treat each variate in a multi-variate time series independently to standardise generative tasks like forecasting, while Goswami et al.
 (2024) extend uni-variate analysis to discriminative tasks like classification.
-Similarly, Jiang et al.
+Jiang et al.
 (2024) and Yang et al.
 (2024) standardise time series by cropping them into segments of predefined size and resampling them to a uniform frequency, respectively, to enable general classification capabilities in medical domains.
-While partially addressing time series heterogeneity, these methods limit model capabilities for general time series analysis.
-Standardisation techniques like cropping or resampling may distort inter-variate relationships, temporal dynamics, and long-range dependencies.
-Additionally, many of these approaches are tailored to specific applications, such as generative tasks (Das et al., 2024; Liu et al., 2024; Woo et al., 2024), or focused on particular domains like medicine (Jiang et al., 2024; Yang et al., 2024).
-Moreover, recent foundational models (Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) focus on uni-variate analysis, ignoring crucial inter-variate relationships essential for real-world applications, such as disease prediction (Schoffelen & Gross, 2009; Wu et al., 2022b).
-Our study aims to overcome these limitations by fully addressing heterogeneity of multi-domain time series, establishing a foundation for general time series analysis across domains and tasks.
+While partially addressing time series heterogeneity, these pre-training paradigms cannot unlock foundational model capabilities.
+Standardisation techniques like cropping or resampling may distort inter-variate relationships and temporal patterns.
+Additionally, many of these approaches are tailored to specific applications, such as generative tasks (Das et al., 2024; Liu et al., 2024; Woo et al., 2024), or focus on domains in a particular field, like medicine (Jiang et al., 2024; Yang et al., 2024).
+Moreover, recent foundation models (Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) focus on uni-variate analysis, ignoring inter-variate relationships that are essential for real-world applications, such as disease prediction (Schoffelen & Gross, 2009; Wu et al., 2022b).
+Our study aims to overcome these limitations by effectively addressing time series heterogeneity, establishing a robust foundation for general time series analysis.
 2.2 Time Series Tokenisation
-Transformers (Vaswani et al., 2017) have emerged as the preferred architecture for foundational models in NLP and CV due to their scalability (Kaplan et al., 2020; Gordon et al., 2021; Alabdulmohsin et al., 2022), enabling the training of models in the magnitude of $100$ billion parameters (Chowdhery et al., 2023; Touvron et al., 2023; Oquab et al., 2024; Ravi et al., 2024).
+Transformers (Vaswani et al., 2017) have emerged as the preferred architecture for foundation models in NLP and CV due to their scalability (Kaplan et al., 2020; Gordon et al., 2021; Alabdulmohsin et al., 2022), enabling the training of large models (Chowdhery et al., 2023; Touvron et al., 2023; Oquab et al., 2024; Ravi et al., 2024).
 To utilise a Transformer for time series analysis, a tokeniser is required to map the time series into a compact latent space.
@@
 While uni-variate models (Nie et al., 2023; Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) consider only temporal positions, multi-variate approaches (Woo et al., 2024; Yang et al., 2024; Jiang et al., 2024) account for both temporal and variate positions.
-However, none of these methods address the unique characteristics of variates, mistakenly assuming that the relationships between variates are identical across domains.
-Our work seeks to adapt the tokenisation process to preserve the domain-specific relationships between variates.
+However, none of the current methods consider the varying characteristics across domains, e.g.
+mistakenly assuming that the relationships between variates are identical across domains.
+Our work seeks to adapt time series tokenisation to account for these domain-specific characteristics.
+Figure 2: Pre-training of OTiS.
+A time series is split into fixed-size patches, which are then embedded using a universal patch projector.
+In addition to a temporal embedding, each patch embedding is modulated with a domain-specific variate embedding to account for the unique characteristics of a domain.
+The resulting input tokens are masked with a dual masking strategy and reconstructed to optimise OTiS.
+The reconstruction is guided by a combination of mean squared error (MSE) and normalised cross-correlation (NCC) loss terms.
 3 Methods
-Figure 2: Architecture of OTiS.
-During pre-training, batches of time series from diverse domains are processed using a domain-specific tokeniser.
-This tokeniser splits a time series into fixed-size patches, which are then embedded using a patch projector shared across all variates and domains.
-A temporal embedding and a domain-specific variate embedding are added to each patch embedding.
-A dual masking strategy is employed to mask the resulting input tokens.
-The reconstruction of the multi-domain input tokens is guided using an auxiliary normalised cross-correlation (NCC) loss.
-In this work, we present a novel multi-domain pre-training paradigm that enables generalisable feature extraction from large, heterogeneous time series corpora.
-We introduce a domain-specific tokeniser with learnable signatures to address heterogeneity in multi-domain time series, as described in Section 3.1.
-We tailor masked data modelling (MDM) for multi-domain time series to pre-train our open model for general time series analysis (OTiS) on a large, heterogeneous corpus, as detailed in Section 3.2.
-In particular, we propose normalised cross-correlation as a loss term to capture global temporal dynamics in time series, as explained in Section 3.3.
-Moreover, we introduce a dual masking strategy to capture bidirectional relationships and temporal causality, essential for general time series analysis, as described in Section 3.4.
-After pre-training, we fine-tune OTiS on limited data to perform a variety of downstream tasks in any - including previously unseen - domain, as outlined in Section 3.5.
-A graphical visualisation of our method is provided in Figure 2.
+In this work, we present a novel multi-domain pre-training paradigm that enables the extraction of generalisable time series features.
+We introduce a tokeniser that accounts for varying time series characteristics across domains (see Section 3.1).
+We tailor masked data modelling to multi-domain time series, enabling our open model for general time series analysis (OTiS) to be pre-trained on a large, heterogeneous corpus (see Section 3.2).
+In particular, we introduce normalised cross-correlation as a loss term to capture global temporal patterns in time series (see Section 3.3) and propose a dual masking strategy to model both bidirectional relationships and temporal causality in time series (see Section 3.4).
+After pre-training, OTiS can be fine-tuned on limited data from any domain, including previously unseen ones, to perform general time series tasks, such as classification, regression, and forecasting (see Section 3.5).
 3.1 Domain-Specific Tokeniser
@@
 We randomly crop or zero-pad $\bm{X}$ to a fixed context length of $\overline{T}$ time points.
-We then split it into $T^{\prime}$ temporal patches of size $P$ along the time dimension, resulting in $V_{S}\cdot T^{\prime}$ patches ${\bm{x}}_{v,t}\in\mathbb{R}^{1\times P}$, where $v\in\{1,\dots,V_{S}\}$ and $t\in\{1,\dots,T^{\prime}\}$.
-Next, we embed these patches using a shared patch projector across all variates and domains, resulting in patch embeddings $e^{\mathcal{P}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}\in\mathbb{R}^{1\times D}$, where $D$ denotes the model dimension.
+We then split it into $T^{\prime}$ temporal patches of size $P$ along the time dimension, resulting in $V_{S}\cdot T^{\prime}$ patches $x_{v,t}\in\mathbb{R}^{1\times P}$, where $v\in\{1,\dots,V_{S}\}$ and $t\in\{1,\dots,T^{\prime}\}$.
+Next, we embed these patches using a universal patch projector shared across domains, producing patch embeddings $e^{\mathcal{P}}(x_{v,t})=e^{\mathcal{P}}_{v,t}\in\mathbb{R}^{1\times D}$, where $D$ denotes the model dimension.
 The patch projector consists of a $1$D convolutional layer followed by layer normalisation and GELU activation.
-The permutation-equivariant nature of Transformers (Vaswani et al., 2017) requires the use of positional embeddings to accurately capture the inherent relationships in the input data.
-Initially introduced for $1$D textual token sequences (Vaswani et al., 2017), positional embeddings simply introduce an ordering into the input sequence.
-Modern implementations further extend their capabilities to encode more complex geometric information, such as $2$D spatial (Dosovitskiy et al., 2021) or graph (Kreuzer et al., 2021) structures.
-For the analysis of any-variate time series, we distinguish between the temporal and variate structure.
-The temporal structure is equivalent to a sequential $1$D structure, such that we use standard $1$D sinusoidal embeddings $e^{\mathcal{T}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{T}}_{t}\in\mathbb{R}^{1\times D}$.
+With a unified patch size and a domain-agnostic design, it serves as a general feature extractor and is inherently frequency-agnostic.
+While the sampling frequency varies across domains, it consistently ensures that relevant information is captured within the observation period, i.e.
+time series.
+Hence, as long as a model’s context length is sufficiently long, it has access to the relevant information within a time series, regardless of the sampling frequency.
+The permutation-equivariant nature of Transformers (Vaswani et al., 2017) requires the use of positional embeddings to accurately capture the inherent relationships of the input.
+Initially introduced for $1$D textual sequences (Vaswani et al., 2017), positional embeddings simply introduce an ordering into the input sequence.
+Modern implementations extend their capabilities to encode more complex geometric information, such as $2$D spatial (Dosovitskiy et al., 2021) or graph (Kreuzer et al., 2021) structures.
+In this work, we differentiate between temporal and variate structures for the analysis of any-variate time series.
+The temporal structure is equivalent to a $1$D sequence, such that we use standard $1$D sinusoidal embeddings $e^{\mathcal{T}}(x_{v,t})=e^{\mathcal{T}}_{t}\in\mathbb{R}^{1\times D}$.
 The variate structure exhibits great heterogeneity across domains.
 In domains with uni-variate and two-variate data, such as mono and stereo audio, the structure is either trivial or only requires a basic distinction between variates.
-In other domains, however, the variate structure may represent more complex relationships, such as $3$D manifolds for electroencephalography (EEG) or electrocardiography (ECG) data, or be of non-spatial nature, such as for financial data.
-Hence, we introduce learnable domain-specific variate embeddings to adequately address the heterogeneity across domains.
-These embeddings, denoted as $e^{\mathcal{V}}_{\mathcal{S}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$ for each variate $v$ in domain $S$, are designed to model the unique properties of a domain.
-They capture the inter-variate relationships and temporal dynamics specific to domain $S$, forming what can be considered as the signature of the very domain.
-Finally, the patch, temporal, and domain-specific variate embeddings are summed to form the input token ${\bm{e}}_{v,t}=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}+\displaystyle{\bm{e}}^{\mathcal{T}}_{t}+\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$.
-These input tokens collectively constitute the final input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
-To support batches of any-variate time series from multiple domains, we pad the variate dimension to the maximum number of variates in a batch $\overline{V}=\max_{S}V_{S}$.
+In other domains, however, the variate structure may represent more complex relationships, such as $3$D manifolds for electroencephalography (EEG) or electrocardiography (ECG), or be of non-spatial nature, e.g.
+for financial data.
+To account for these varying characteristics across domains, we introduce learnable domain-specific variate embeddings.
+These embeddings, denoted as $e^{\mathcal{V}}_{S}(x_{v,t})=e^{\mathcal{V}}_{S,v}\in\mathbb{R}^{1\times D}$ for each variate $v$ in domain $S$, are designed to capture the unique characteristics of a domain $S$, including the domain-specific inter-variate relationships and temporal patterns.
+Hence, they can be considered as a domain signature after training.
+Finally, the patch, temporal, and domain-specific variate embeddings are summed to form the input token $e_{v,t}=e^{\mathcal{P}}_{v,t}+e^{\mathcal{T}}_{t}+e^{\mathcal{V}}_{S,v}\in\mathbb{R}^{1\times D}$.
+These tokens collectively form the final input sequence $\bm{E}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
+As time series across domains vary in the number of variates, we pad the variate dimension to the maximum number of variates $\overline{V}=\max_{S}V_{S}$ within each batch.
 For samples where $V_{S}<\overline{V}$ or $T<\overline{T}$, attention masking is used to ensure that padded variate or temporal tokens are ignored.
 The domain-specific tokeniser is trained end-to-end with the Transformer layers.
-Definition of (Sub-)Domains.
+Definition of Domains.
 The domain-specific tokeniser is designed to integrate different datasets within a domain.
@@
 In this case, a single EEG-specific tokeniser ($V_{\text{EEG}}=19$) is sufficient to accommodate both sampling frequencies, i.e.
-${\bm{E}}^{\mathcal{V}}_{\text{EEG-TDBrain}}=\displaystyle{\bm{E}}^{\mathcal{V}}_{\text{EEG-SEED}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},19}]^{\top}\in\mathbb{R}^{19\times D}$, as demonstrated in our experiments in Section 4.
-Note that while these positional embeddings are agnostic to variate ordering, we simplify processing by aligning the variate order across datasets within the same domain.
-Consider another EEG dataset, LEMON (Babayan et al., 2019), which includes $62$ electrodes.
-Of these, $15$ overlap with the electrodes in TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), while the remaining $47$ are unique to LEMON (Babayan et al., 2019).
-In this scenario, the EEG-specific tokeniser can be extended by the $47$ new variates ($V_{\text{EEG}}=66$), such that ${\bm{E}}^{\mathcal{V}}_{\text{EEG-LEMON}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},15},\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},20},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},66}]^{\top}\in\mathbb{R}^{62\times D}$.
-In this way, different datasets can be combined to approximate the underlying data distribution of a domain $S$, e.g.
+$\bm{E}^{\mathcal{V}}_{\text{EEG-TDBrain}}=\bm{E}^{\mathcal{V}}_{\text{EEG-SEED}}=[e^{\mathcal{V}}_{\text{EEG},1},\dots,e^{\mathcal{V}}_{\text{EEG},19}]^{\top}\in\mathbb{R}^{19\times D}$, as shown in our experiments in Section 4.
+Note that while these positional embeddings are agnostic to variate ordering, we simplify processing by aligning the variate order across datasets within a domain.
+Consider another EEG dataset, LEMON (Babayan et al., 2019), with $62$ variates.
+Of these, $15$ overlap with the variates in TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), while the remaining $47$ are unique to LEMON (Babayan et al., 2019).
+In this scenario, the EEG-specific tokeniser can be extended by the $47$ new variates ($V_{\text{EEG}}=66$), such that $\bm{E}^{\mathcal{V}}_{\text{EEG-LEMON}}=[e^{\mathcal{V}}_{\text{EEG},1},\dots,e^{\mathcal{V}}_{\text{EEG},15},e^{\mathcal{V}}_{\text{EEG},20},\dots,e^{\mathcal{V}}_{\text{EEG},66}]^{\top}\in\mathbb{R}^{62\times D}$.
+In this way, different datasets can be combined to approximate the underlying data distribution of a domain, e.g.
 EEG, enabling the creation of large and diverse time series corpora.
@@
 3.2 Pre-Training on Multi-Domain Time Series
-We pre-train our model using masked data modelling (MDM) (He et al., 2022) to learn generalisable time series features across domains.
-We mask a subset of input tokens and only encode the visible (i.e.
-non-masked) tokens using an encoder $f(\cdot)$.
-Afterwards, we complement the encoded tokens with learnable mask tokens and feed them to a decoder $g(\cdot)$, reconstructing the original input tokens.
-More precisely, we draw a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$, following the dual masking strategy proposed in Section 3.4, and apply it to the input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
-Thus, we obtain a visible view ${\bm{E}}[\displaystyle{\bm{m}}]\in\mathbb{R}^{N_{1}\times D}$, where $N_{1}=\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}m_{v,t}$ and $N_{0}=(V_{S}\cdot T^{\prime})-N_{1}$ denote the number of visible and masked tokens, respectively.
-The visible view ${\bm{E}}[\displaystyle{\bm{m}}]$ is then fed to the encoder $f(\cdot)$ to compute the token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$:
-${\bm{H}}=f(\displaystyle{\bm{E}}[\displaystyle{\bm{m}}]).$ (1)
-To reconstruct the original input, these token features are fed to the decoder $g(\cdot)$ together with a special, learnable mask token ${\bm{e}}^{\mathcal{M}}\in\mathbb{R}^{1\times D}$, that is inserted at the masked positions where $m_{v,t}=0$:
-${\bm{h}}^{\prime}_{v,t}=\begin{cases}\displaystyle{\bm{h}}_{v,t}&\text{if }m_{v,t}=1\\ \displaystyle{\bm{e}}^{\mathcal{M}}&\text{if }m_{v,t}=0\end{cases},$ (2)
+To learn generalisable time series features, we adapt masked data modelling (MDM) (He et al., 2022) and pre-train OTiS on multi-domain time series, as visualised in Figure 2.
+We mask a subset of the input tokens and only encode the visible tokens using an encoder $f(\cdot)$.
+We then complement the encoded tokens with learnable mask tokens and pass them through a decoder $g(\cdot)$ to reconstruct the input tokens.
+More precisely, we draw a binary mask $m\in\{0,1\}^{V_{S}\cdot T^{\prime}}$, following a dual masking strategy introduced in Section 3.4, and apply it to the input sequence $\bm{E}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
+Thus, we obtain a visible view $\bm{E}[m]\in\mathbb{R}^{N_{1}\times D}$, where $N_{1}=\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}m_{v,t}$ and $N_{0}=(V_{S}\cdot T^{\prime})-N_{1}$ denote the number of visible and masked tokens, respectively.
+The visible tokens $E[m]$ are then fed to the encoder $f(\cdot)$ to compute the time series features $\bm{H}\in\mathbb{R}^{N_{1}\times D}$:
+$\bm{H}=f(\bm{E}[m]).$ (1)
+To reconstruct the original input tokens, these features are passed through the decoder $g(\cdot)$ along with a learnable mask embedding $e^{\mathcal{M}}\in\mathbb{R}^{1\times D}$, which is inserted at each masked position where $m_{v,t}=0$:
+$h^{\prime}_{v,t}=\begin{cases}h_{v,t}&\text{if }m_{v,t}=1\\ e^{\mathcal{M}}+e^{\mathcal{T}}_{t}+e^{\mathcal{V}}_{S,v}&\text{if }m_{v,t}=0\end{cases},$ (2)
 such that $\bm{H}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
-The decoder $g(\cdot)$ then predicts the reconstructed input $\widehat{{\bm{X}}}\in\mathbb{R}^{V_{S}\times(T^{\prime}\cdot P)}$:
-$\widehat{{\bm{X}}}=g(\displaystyle{\bm{H}}^{\prime}),$ (3)
+The decoder $g(\cdot)$ then predicts the reconstructed input $\bm{\widehat{X}}\in\mathbb{R}^{V_{S}\times(T^{\prime}\cdot P)}$:
+$\bm{\widehat{X}}=g(\bm{H}^{\prime}),$ (3)
 where $(T^{\prime}\cdot P)=\overline{T}$, i.e.
@@
 Eventually, the domain-specific tokeniser described in Section 3.1, the encoder $f(\cdot)$, and the decoder $g(\cdot)$ are optimised end-to-end using the mean squared error (MSE) loss on all reconstructed input tokens:
-$\mathcal{L}_{\text{MSE}}=\frac{1}{V_{S}\cdot T^{\prime}}\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}\|\displaystyle{\bm{x}}_{v,t}-\displaystyle\widehat{{\bm{x}}}_{v,t}\|_{2}^{2}.$ (4)
+$\mathcal{L}_{\text{MSE}}=\frac{1}{V_{S}\cdot T^{\prime}}\sum_{v=1}^{V_{S}}\sum_{t=1}^{T^{\prime}}\|x_{v,t}-\widehat{x}_{v,t}\|_{2}^{2}.$ (4)
 3.3 Normalised Cross-Correlation Loss
@@
 However, time series often exhibit long-range dependencies, where past values influence future outcomes over extended periods.
-To accurately capture these global patterns, we introduce normalised cross-correlation (NCC) as a loss term in MDM for time series:
-$\mathcal{L}_{\text{NCC}}=\frac{1}{V_{S}\cdot\overline{T}}\sum^{V_{S}}_{v=1}\sum^{\overline{T}}_{t=1}\hskip 2.84526pt\frac{1}{\sigma_{\displaystyle{\bm{x}}_{v}}\sigma_{\displaystyle\widehat{{\bm{x}}}_{v}}}(x_{v,t}-\mu_{\displaystyle{\bm{x}}_{v}})(\widehat{x}_{v,t}-\mu_{\displaystyle\widehat{{\bm{x}}}_{v}})\in\left[-1,1\right],$ (5)
-where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively. Hence, to capture both local and global temporal dynamics, the total loss used to optimise OTiS is defined as
+To accurately capture these global patterns, we introduce normalised cross-correlation (NCC) as a loss term
+$\mathcal{L}_{\text{NCC}}=\frac{1}{V_{S}\cdot\overline{T}}\sum^{V_{S}}_{v=1}\sum^{\overline{T}}_{t=1}\hskip 2.84526pt\frac{1}{\sigma_{x_{v}}\sigma_{\widehat{x}_{v}}}(x_{v,t}-\mu_{x_{v}})(\widehat{x}_{v,t}-\mu_{\widehat{x}_{v}})$ (5)
+for MDM in time series, where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively, and $\mathcal{L}_{\text{NCC}}\in\left[-1,1\right]$.
+Hence, to capture both local and global temporal patterns, the total loss used to optimise OTiS is defined as
 $\mathcal{L}=\mathcal{L}_{\text{MSE}}+\lambda\cdot\left(1-\mathcal{L}_{\text{NCC}}\right),$ (6)
-where $\lambda$ is empirically set to $0.1$ during pre-training.
+where $\lambda\in\mathbb{R}$ is empirically set to $0.1$ during pre-training.
 3.4 Dual Masking Strategy
@@
 Specifically, we randomly select between two masking schemes during pre-training, namely random masking and post-fix masking.
-In 757575\,75of cases, we apply random masking, where each $m_{v,t}$ is independently sampled from a Bernoulli distribution with probability $p=1-\rho$, with $\rho$ denoting the masking ratio (i.e.
+In 757575\,75of cases, we apply random masking, where each $m_{v,t}$ is independently sampled from a Bernoulli distribution with probability $p=1-\rho$, with $\rho\in[0,1]$ denoting the masking ratio (i.e.
 $m_{v,t}\sim\text{Bernoulli}(1-\rho)$).
@@
 The prediction of future values solely based on past observations simulates real-world forecasting conditions, helping the model to capture temporal causality.
-Overall, this dual masking strategy enables OTiS to learn both bidirectional relationships and temporal causality, which are essential for general time series analysis.
-3.5 Fine-Tuning & Inference on (Unseen) Target Domains
-Inclusion of Unseen Domains.
-For a new domain $S$, a randomly initialised variate embedding ${\bm{E}}^{\mathcal{V}}_{\mathcal{S}}\in\mathbb{R}^{V_{S}\times D}$ is introduced.
-The domain-specific tokeniser is then fine-tuned alongside the encoder $f(\cdot)$, and, if required, the decoder $g(\cdot)$, for the specific downstream task, as described in the following.
+Note that the optimal balance between the two masking schemes is empirically determined (see Appendix G.1).
+Overall, this dual masking strategy enables OTiS to learn both bidirectional relationships and temporal causality, which prove essential for general time series analysis.
+3.5 Fine-Tuning & Inference
+Adaptation to Unseen Domains.
+For a previously unseen domain $S$, we randomly initialise variate embeddings $\bm{E}^{\mathcal{V}}_{S}\in\mathbb{R}^{V_{S}\times D}$.
+These embeddings are then fine-tuned along with the pre-trained encoder $f(\cdot)$ and, if required, the decoder $g(\cdot)$, for the specific task in $S$, as detailed in the following.
 Classification & Regression.
-We use the encoder $f(\cdot)$ and the unmasked input sequence ${\bm{E}}$ to compute all token features ${\bm{H}}=f(\displaystyle{\bm{E}})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
-We average-pool these features into a global token ${\bm{h}}^{*}\in\mathbb{R}^{1\times D}$, which we feed through a linear layer to obtain the final model prediction.
+We use the pre-trained encoder $f(\cdot)$ and the unmasked input sequence $\bm{E}$ to extract the time series features $\bm{H}=f(\bm{E})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
+We average-pool these features into a global representation $h^{*}\in\mathbb{R}^{1\times D}$, which is passed through a linear layer to obtain the final model prediction.
 We optimise a cross-entropy and MSE loss for the classification and regression tasks, respectively.
 Forecasting.
-We apply post-fix masking to generate a binary mask ${\bm{m}}\in\{0,1\}^{V_{S}\cdot T^{\prime}}$ for the forecasting task.
-The encoder $f(\cdot)$ is used to compute the visible token features ${\bm{H}}\in\mathbb{R}^{N_{1}\times D}$.
-We then concatenate the sequence with learnable mask tokens to form ${\bm{H}}^{\prime}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$, which is passed through the decoder $g(\cdot)$ to produce the final output.
-We optimise the MSE loss together with the NCC loss term over all reconstructed input tokens.
+We use the pre-trained encoder $f(\cdot)$ and the unmasked input sequence $\bm{E}$ to extract the time series features $\bm{H}=f(\bm{E})\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$.
+We then concatenate the sequence with learnable mask embeddings $e^{\mathcal{M}}$ to form $\bm{H}^{\prime}\in\mathbb{R}^{(V_{S}\cdot(T^{\prime}+T^{\prime}_{H}))\times D}$, where $T^{\prime}_{H}\cdot P$ denotes the forecasting horizon specified in time points.
+This input sequence is then passed through the pre-trained decoder $g(\cdot)$ to obtain the final model prediction.
+We optimise both the MSE and NCC loss terms over all reconstructed input tokens.
 4 Experiments & Results
 4.1 Model Variants and Implementation Details
-We introduce OTiS in three different configurations, Base, Large, and Huge, with their specific architectures described in Appendix C.1, to explore scaling laws with respect to the model size.
-We set the patch size and stride to $P=24$, respectively, to split the time series into $T^{\prime}=\frac{\overline{T}}{P}$ non-overlapping patches along the time dimension.
-For pre-training, the context length specified in time points is set to $\overline{T}=1008$, resulting in $T^{\prime}=42$ sinusoidal temporal embeddings.
+We introduce OTiS in three different configurations - Base, Large, and Huge (see Appendix C.1) - to explore scaling laws with respect to the model size.
+We set the patch size and stride to $P=24$, respectively, to split the time series into $T^{\prime}=\frac{\overline{T}}{P}$ non-overlapping patches along the temporal dimension.
+For pre-training, the context length is set to $\overline{T}=1008$ time points, resulting in $T^{\prime}=42$ sinusoidal temporal embeddings.
 If longer context lengths are required during fine-tuning, these embeddings are linearly interpolated (i.e.
@@
 An overview of the computational costs is provided in Appendix D.
+Table 1: Overview of our pre-training corpus that spans eight domains from medicine, engineering, natural sciences, and finance.
+Domain Name Samples Variates Time points Frequency
+ECG MIMIC-IV-ECG $400,000$ $12$ $5,000$ 500500500\,500Hz
+Temperature DWD $203,340$ $1$ $720$ $278\,\mu$Hz
+Audio (stereo) AudioSet-$20$K $16,123$ $2$ $441,000$ 44.144.144.1\,44.1kHz
+Audio (mono) AudioSet-$20$K $3,491$ $1$ $441,000$ 44.144.144.1\,44.1kHz
+Electromechanics FD-A $13,640$ $1$ $5,120$ 646464\,64kHz
+EEG TDBrain $2,692$ $19$ $60,000$ 500500500\,500Hz
+EEG SEED $675$ $19$ $37,000$ 200200200\,200Hz
+Banking NN5555 $111$ $1$ $971$ $12\,\mu$Hz
+Economics FRED-MD $107$ $1$ $728$ 386386386\,386nHz
+Economics Exchange $8$ $1$ $7,588$ $12\,\mu$Hz
+640,187 11,052,756,981
 4.2 Large and Diverse Pre-Training Corpus
 We aim to develop a general time series model that fully handles the heterogeneity in real-world data.
-Specifically, our model is designed to handle time series with different variate counts $V_{S}$, inter-variate relationships, temporal dynamics, and sampling frequency, ensuring flexibility for downstream tasks.
+Specifically, our model is designed to handle time series with different variate counts $V_{S}$, inter-variate relationships, temporal patterns, and sampling frequency, ensuring flexibility for downstream tasks.
 To this end, we pre-train our model on a large and diverse corpus of publicly available data spanning $8$ domains, with a total of $640,187$ samples and $11$ billion time points, as summarised in Table 1.
@@
 The time series corpus is split into $612,394$ training and $27,793$ validation samples for pre-training.
-Table 1: Overview of our large and diverse pre-training corpus. The corpus is built with unlabelled data from eight domains, encompassing medicine, engineering, natural sciences, and finance.
-Domain Name Samples Variates Time points Frequency Disk size
-$S$ $V_{S}$
-ECG MIMIC-IV-ECG 2023 $400,000$ $12$ $5,000$ 500500500\,500Hz 909090\,90GB
-Temperature DWD 2024 $203,340$ $1$ $720$ (hourly) 278278278\,278mHz 614614614\,614MB
-Audio (stereo) AudioSet-$20$K 2017 $16,123$ $2$ $441,000$ 44.144.144.1\,44.1kHz 535353\,53GB
-Audio (mono) AudioSet-$20$K 2017 $3,491$ $1$ $441,000$ 44.144.144.1\,44.1kHz 666\,6GB
-Electromechanics FD-A 2016 $13,640$ $1$ $5,120$ 646464\,64kHz 161161161\,161MB
-EEG TDBrain 2022 $2,692$ $19$ $60,000$ 500500500\,500Hz 121212\,12GB
-EEG SEED 2015 $675$ $19$ $37,000$ 200200200\,200Hz 222\,2GB
-Banking NN5555 2012 $111$ $1$ $971$ (daily) $12\,\mu$Hz 370370370\,370KB
-Economics FRED-MD 2016 $107$ $1$ $728$ (monthly) 386386386\,386nHz 330330330\,330KB
-Economics Exchange 2018 $8$ $1$ $7,588$ (daily) $12\,\mu$Hz 240240240\,240KB
-640,187 11,052,756,981 164 GB
 4.3 Benchmarking Across Domains and Tasks
-To assess the utility of OTiS in real-world settings, we conduct experiments on three key use cases in time series analysis: classification, regression, and forecasting.
-For classification, we perform binary epilepsy detection using EEG (Epilepsy 2001), multi-class fault detection in rolling bearings from vibration signals (FD-B 2016), multi-class hand-gesture classification with accelerometer signals (Gesture 2009), and multi-class muscular disease classification using electromyographie (EMG 2000).
-In regression, we predict five imaging-derived cardiac phenotypes from $12$-lead ECG (LVEDV, LVESV, LVSV, LVEF, LVM 2020).
-For forecasting, we predict electricity transformer temperature (ETT 2021), weather (Weather 2024), and electricity consumption (Electricity 2024).
-Further dataset details are provided in Appendix B.
-We adhere to the established data splitting and evaluation procedures for the classification (Zhang et al., 2022), regression (Turgut et al., 2023), and forecasting (Zhou et al., 2021) tasks.
-All experiments are reported as mean and standard deviation across five seeds set during fine-tuning.
-Table 2: Classification and regression performance on a total of $9$ benchmark tasks.
-OTiS is competitive with specialised baselines, setting new state-of-the-art on $6$ tasks and even outperforming the multimodal CM-AE and MMCL.
-This demonstrates the capability of OTiS to extract high-level semantics.
+We evaluate OTiS across general use cases in time series analysis: classification, regression, and forecasting.
+We use $12$ datasets from $8$ domains to compare our model against $27$ baselines, as outlined in Appendix B.
+The baselines comprise $15$ specialised, i.e.
+task-specific, models (either fully supervised or pre-trained and fine-tuned exclusively on target data) and $12$ general models (pre-trained on external data and fine-tuned on target data), including $6$ foundation models.
+We follow established data splitting and evaluation procedures for classification (Zhang et al., 2022), regression (Turgut et al., 2025), and forecasting (Zhou et al., 2021), with results reported across five seeds set during fine-tuning.
+Table 2: Classification and regression results on a total of $9$ benchmark tasks, with indicating tasks in previously unseen domains.
 Best score in bold, second best underlined.
-indicates tasks in previously unseen domains.
+OTiS successfully captures high-level semantics for classification and localised time series features for regression.
+Zero-shot and linear probing predictions indicate the generalisability of OTiS’ time series features.
 (a) Classification [Accuracy (ACC1$\uparrow$) in %]
@@
 SimCLR 2020 90.71 49.17 48.04 61.46
+TimesNet 2022a 94.01 56.86 59.79 91.22
 CoST 2022 88.40 47.06 68.33 53.65
@@
 OTiS-Huge 91.48 98.32 63.61 98.37
+OTiS0 95.18 61.32 51.67 95.12
+Zero-shot predictions of OTiS-Base.
 (b) Regression [R-squared ($R^{2}$uparrow)]
 Model LVEDV LVESV LVSV LVEF LVM
-ViT 2023 0.409 0.396 0.299 0.175 0.469
-MAE 2023 0.486 0.482 0.359 0.237 0.573
+ViT 2021 0.409 0.396 0.299 0.175 0.469
+MAE 2022 0.486 0.482 0.359 0.237 0.573
+iTransformer 2023 0.307 0.279 0.227 0.070 0.361
 CM-AE* 2023 0.451 0.380 0.316 0.103 0.536
-MMCL* 2023 0.504 0.503 0.370 0.250 0.608
+MMCL* 2025 0.504 0.503 0.370 0.250 0.608
 OTiS-Base 0.509 0.512 0.391 0.292 0.592
@@
 OTiS-Huge 0.505 0.510 0.376 0.281 0.593
-* Models that incorporate paired imaging data during pre-training.
-OTiS demonstrates strong classification capabilities, setting two new benchmarks, as shown in Table 2(a).
-Additionally, OTiS excels in predicting imaging-derived cardiac phenotypes, even surpassing multimodal baselines that incorporate imaging data during pre-training, as summarised in Table 2(b).
-Our model outperforms the baselines in $4$ out of $5$ regression tasks.
-As shown in Table 3, OTiS also exhibits strong forecasting capabilities, outperforming current models in $4$ out of $6$ benchmarks.
-Notably, all of the forecasting tasks are performed in previously unseen domains.
-A visualisation of the forecast predictions can be found in Appendix F.
-Overall, OTiS outperforms both specialised and general state-of-the-art baselines in $10$ out of $15$ diverse applications across $8$ domains, demonstrating its strong utility and generalisability across (unseen) domains and tasks.
-Table 3: Forecasting performance on $6$ benchmark tasks.
-OTiS is competitive with specialised and general baselines, setting new state-of-the-art on $4$ tasks and showcasing its ability to capture local time series features.
+[…]
+* Models incorporate paired imaging data during pre-training.
+Linear probing of OTiS-Base.
+Table 3: Forecasting results on $6$ benchmark tasks, with indicating tasks in previously unseen domains.
 A forecasting horizon of $96$ time points is predicted from the past $336$ (*$512$, ${}^{+}904$) time points.
@@
 Best score in bold, second best underlined.
-indicates tasks in previously unseen domains.
+OTiS successfully captures temporal patterns and trends essential for forecasting.
+Predictions involving minimal training () indicate the generalisability of OTiS’ extracted time series features.
 Model ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
@@
 OTiS-Huge 0.461 0.215 0.384 0.181 0.149 0.132
+[…]
 Experiments could not be conducted on a single NVIDIA RTX A6000600060006000-$48$GB GPU.
+Model incorporates paired text data during pre-training and fine-tuning.
+Predictions of OTiS-Base with only the domain-specific variate embeddings (VE) being fine-tuned.
+The experiments reveal that OTiS extracts powerful time series features, achieving state-of-the-art performance on $10$ out of $15$ diverse benchmarks.
+The classification results in Table 2(a) highlight its particular strength in processing long time series, yielding a huge performance boost on FD-B ($\overline{T}_{\text{FD-B}}=5112$).
+The regression results in Table 2(b) emphasise the effectiveness of pre-training across domains, with OTiS outperforming specialised models that are unimodally pre-trained (e.g.
+MAE) or even multimodally pre-trained (e.g.
+MMCL).
+This is further supported by an additional ablation study on different training strategies presented in Appendix G.2.
+Moreover, the forecasting results in Table 3 and Appendix H reveal that OTiS effectively captures temporal patterns and trends, even in previously unseen domains.
+Finally, zero-shot experiments detailed in Appendix F, along with experiments involving minimal training (e.g.
+linear probing or fine-tuning of domain-specific embeddings) in Table 2(b) and 3, highlight how the extracted time series features generalise across domains and tasks, e.g.
+evidenced by the Epilepsy, LVEDV, and ETTh2 benchmarks.
+(a) Frequency.
+(b) Amplitude.
+(c) Offset.
+(d) Phase.
+(e) Composition.
+Figure 3: First two principal components of zero-shot sinusoidal representations extracted by OTiS-Base.
+We freeze OTiS after pre-training and use randomly initialised variate embeddings.
+The output features of the encoder are averaged to obtain a global representation.
+OTiS has an intrinsic understanding of fundamental time series properties, such as (a) frequency, (b) amplitude, (c) offset, and (d) phase.
+It effectively disentangles (e) simultaneous variations of these properties, providing a strong foundation for general time series analysis.
+4.4 General Time Series Understanding
+To evaluate whether OTiS has an intrinsic understanding of fundamental time series properties, such as frequency, amplitude, offset, and phase, we analyse zero-shot representations of sine waves, as illustrated in Figure 3.
+Particularly, we evaluate OTiS after pre-training without fine-tuning, using randomly initialised variate embeddings.
+We average the encoder’s output features to obtain a global representation.
+When examining time series properties in isolation, we observe that their continuous variations are correctly mirrored in OTiS’ latent space.
+For frequency (see Figure 3(a)), the analysis reveals that embeddings for higher frequencies are positioned closer together, indicating greater similarity, than those for lower frequencies.
+This observation aligns with intuition: a change in frequency from 111\,1Hz to 505050\,50Hz represents a factor of $50$, thus resulting in more distinct embeddings, whereas a change from 505050\,50Hz to 100100100\,100Hz reflects only a doubling of the frequency.
+Analogous patterns are observed for amplitude (see Figure 3(b)) and offset (see Figure 3(c)).
+In the case of phase (see Figure 3(d)), the latent space forms a closed circle, with embeddings for phases of 00 and $2\pi$ being identical and positioned opposite to the embedding for a phase of $\pi$.
+These findings are consistent with the mathematical definition of a unit circle in the complex plane:
+$e^{i\phi}=\text{cos}(\phi)+i\hskip 1.0pt\text{sin}(\phi),$ (7)
+where $\phi\in[0,2\pi]$ denotes the phase.
+A phase of $\phi=2\pi$ corresponds to the identity (i.e.
+$e^{i2\pi}=e^{i0}$), while $\phi=\pi$ corresponds to a multiplication by $-1$ (i.e.
+$e^{i\pi}=-e^{i0}$).
+Moreover, when faced with simultaneous variations in multiple time series properties (see Figure 3(e)), we observe that our model successfully disentangles them in its latent space.
+Overall, these findings demonstrate an intrinsic understanding of fundamental time series properties, providing a strong and essential foundation for general time series analysis.
 4.5 Domain Signature Analysis
-A key component of OTiS is its use of domain-specific variate embeddings.
-While these embeddings are randomly initialised, we expect them to capture unique domain characteristics during training, eventually serving as the signature of their respective domain.
-To validate this hypothesis, we analyse the domain-specific variate embeddings after pre-training using principal component analysis (PCA).
-First, we find that OTiS unifies time series from diverse domains into a meaningful latent space, where embeddings of domains with shared high-level semantics cluster together, as depicted in Appendix E.1.
-For example, embeddings of mono and stereo audio group closely, as do those of banking and economics.
-Moreover, EEG-specific embeddings are clearly separated and ECG-specific embeddings form a tight cluster.
-Second, OTiS preserves low-level semantics specific to each domain, such as the relationships between variates.
-To explore this, we focus on the learned variate embeddings of EEG, the most complex domain in our corpus.
-EEG variates correspond to actual electrodes, each associated with a $3$D position in space or a $2$D position on the scalp, which is ideal for studying inter-variate relationships.
-Our analysis covers both (i) variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during multi-domain pre-training, and (ii) variate embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning of the pre-trained OTiS.
-We find that the first three principal components explain (i) $74.7\,\%$ and (ii) $87.9\,\%$ of the variance, indicating that the $3$D PCA projections of the variate embeddings correspond closely to the actual electrode positions.
-A visualisation of the learned $3$D electrode layout is provided in Appendix E.
-Additionally, we linearly align these positions with the true EEG electrode layout on the $2$D scalp manifold, as depicted in Figure 3.
-The alignment yields $R^{2}$ values of (i) $0.81$ and (ii) $0.95$, confirming a strong correspondence between the learned variate embeddings and the actual electrode layout.
-Figure 3: $2$D PCA projections of the learned EEG-specific variate embeddings, overlaid on the actual EEG electrode layout.
-(Left) Embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training.
-(Right) Embeddings for previously unseen EEG recordings with $32$ electrodes learned during fine-tuning.
-The embeddings accurately reflect the spatial electrode layout, as confirmed by high correlations (R2) between the PCA projections $\bullet$ and the true layout $\circ$.
+Inter-Variate Relationships.
+A key design choice of OTiS is the use of domain-specific variate embeddings.
+While these embeddings are randomly initialised, we expect them to capture unique domain characteristics, e.g.
+inter-variate relationships, during training.
+To test this hypothesis, we analyse the EEG-specific and ECG-specific variate embeddings learned during pre-training, and the Weather-specific variate embeddings learned during fine-tuning.
+The principal component analysis (PCA) presented in Figure 7 (see Appendix E.2.1) demonstrates that EEG-specific variate embeddings accurately capture the $3$D spatial arrangement of EEG electrodes placed on the human scalp.
+Similarly, the PCA in Figure 9 (see Appendix E.2.1) indicates that ECG-specific variate embeddings correctly reflect the relationships among the $12$ leads: limb leads I, II, and III form a planar $2$D triangle, representing Einthoven’s triangle (Kligfield et al., 2007), while precordial leads V1111–V6666 capture a $3$D view of the heart.
+Moreover, the embedding similarity analysis in Figure 10 (see Appendix E.2.1) reveals that Weather-specific variate embeddings capture the true physical relationships among climatological indicators.
+This also showcases that OTiS successfully adapts to previously unseen domains with limited data seen during fine-tuning.
+Temporal Patterns.
+The central question is whether OTiS learns only the relationships across variates or also captures temporal patterns within a variate.
+To investigate this, we conduct forecasting experiments on uni-variate sine waves with distinct frequencies.
+Note that in a uni-variate setting ($V_{\text{sine}}=1$), we ensure that our model does not utilise information from other variates.
+For these experiments, we freeze both the encoder $f(\cdot)$ and decoder $g(\cdot)$ after pre-training and fine-tune only a randomly initialised variate embedding $\bm{E}^{\mathcal{V}}_{\text{sine}}\in\mathbb{R}^{1\times D}$ ($<0.2\,$k trainable parameters) on 505050\,50Hz sine waves.
+During inference, OTiS is fully frozen and forecasts are made on unseen sine waves with frequencies of 222\,2Hz, 282828\,28Hz, 606060\,60Hz, and 100100100\,100Hz.
+As shown in Figure 11 (see Appendix E.2.2), we find that the domain-specific variate embedding $\bm{E}^{\mathcal{V}}_{\text{sine}}$ captures temporal patterns that generalise to previously unseen frequencies.
+These observations demonstrate that OTiS learns not only the inter-variate relationships of a domain, but also its temporal patterns.
 4.6 Scaling Study
-Figure 4: Performance of OTiS with different numbers of pre-training samples.
+Figure 4: Scaling study.
 Shaded regions indicate the standard deviation across $5$ seeds.
-Increasing dataset size generally improves downstream performance.
-Scaling model size requires even larger pre-training corpora to be effective.
-We analyse the scaling behaviour of OTiS with respect to model and dataset size.
+Downstream performance across tasks generally scales with dataset size.
+Scaling the model size requires even larger pre-training corpora to be effective.
+We analyse the scaling behaviour of OTiS with respect to dataset and model size.
 To this end, we subsample the pre-training data to $10\,\%$ and $1\,\%$ of its original size, ensuring that each subset is fully contained within the corresponding superset.
@@
 The experiments demonstrate that downstream performance generally scales with dataset size, achieving the best results with the full pre-training dataset.
-This trend, however, does not directly apply to model size, which is in line with the scaling behaviour observed in current time series foundational models (Woo et al., 2024; Goswami et al., 2024).
-Given that performance generally improves across all models with increasing data size, we hypothesise that scaling the model size could prove beneficial with even larger pre-training corpora.
+This trend, however, does not directly apply to model size, which is in line with the scaling behaviour observed in current time series foundation models (Woo et al., 2024; Goswami et al., 2024).
+Given that an increasing dataset size generally improves model performance, we hypothesise that scaling the model size could prove beneficial with even larger pre-training corpora.
 4.7 Ablation Study
-Figure 5: Ablation study on key components of OTiS.
+We conduct an ablation study to analyse the impact of OTiS’ key components: the domain-specific tokeniser, dual masking strategy, and normalised cross-correlation (NCC) loss.
+As shown in Figure 5, OTiS achieves superior performance when all components are integrated during pre-training.
+Replacing domain-specific variate embeddings with domain-agnostic ones (i.e.
+variate embeddings shared across all domains) consistently led to inferior performance across tasks, indicating the importance of considering domain characteristics during tokenisation.
+Switching from dual to random masking resulted in performance degradation, although the impact was less notable for generative tasks than for discriminative tasks.
+We hypothesise that the NCC loss already captures temporal causality, which is particularly helpful for generative tasks like forecasting.
+Overall, removing the NCC loss caused performance declines across all downstream applications, emphasising the importance of long-range dependencies for general time series understanding.
+Figure 5: Ablation study.
 Downstream performance is analysed across $5$ seeds.
 A leave-one-out approach is used to evaluate the influence of each component.
-The default setting, which includes all components, demonstrates superior model capabilities across tasks.
-We perform an ablation study to analyse the impact of OTiS’ key components: the domain-specific tokeniser, dual masking strategy, and normalised cross-correlation (NCC) loss.
-As shown in Figure 5, the best and most robust performance is achieved when all components are used during pre-training.
-Replacing the domain-specific variate embeddings with domain-agnostic embeddings (i.e.
-learnable embeddings shared across all domains) consistently led to inferior performance across all tasks, demonstrating the importance of capturing domain-specific data characteristics during tokenisation.
-Switching from dual masking to random masking resulted in performance degradation, although the impact was less notable for generative tasks than for discriminative tasks.
-We hypothesise that the NCC loss already captures temporal causality, which is particularly crucial for generative tasks like forecasting.
-Overall, removing the NCC loss caused performance declines across all downstream tasks, emphasising the role of long-range dependencies for general time series understanding.
+The default setup, incorporating all components, achieves superior performance across tasks.
 5 Discussion & Conclusion
-In this study, we explore the problem of effective pre-training on heterogeneous time series corpora.
-Time series vary substantially across domains, e.g.
-with respect to inter-variate relationships and temporal dynamics, rendering generalisable feature extraction from multi-domain time series difficult.
-To address this issue, we present OTiS, an open model for general time series analysis, specifically designed to handle multi-domain heterogeneity.
-Our novel multi-domain pre-training paradigm, including a domain-specific tokeniser with learnable signatures, a dual masking strategy, and a normalised cross-correlation (NCC) loss, enables OTiS to extract generalisable time series features.
-In extensive experiments, we demonstrate that OTiS generalises well across $15$ diverse downstream applications spanning $8$ distinct domains, achieving competitive performance with both specialised and general state-of-the-art (SOTA) models.
-In a qualitative analysis, we further show that OTiS unifies time series from diverse domains in a meaningful latent space, while preserving low-level semantics of a domain including the inter-variate relationships.
-Thereby, our work establishes a strong foundation for future advancements in interpretable and general time series analysis.
+In this study, we explore the problem of efficient pre-training on heterogeneous time series corpora.
+Prevalent in domains ranging from medicine to finance, time series vary notably e.g with respect to inter-variate relationships and temporal patterns, rendering efficient pre-training across domains difficult.
+To address this, we propose a novel pre-training paradigm including a tokeniser with learnable domain signatures, a dual masking strategy, and a normalised cross-correlation loss, enabling our open model for general time series analysis (OTiS) to learn general time series features.
+In extensive benchmarking experiments, OTiS is competitive against specialised state-of-the-art (SoTA) models.
+Our qualitative analysis highlights how learnable domain signatures address time series heterogeneity, facilitating the disentanglement of domain-specific knowledge and enabling OTiS to acquire knowledge on fundamental time series properties shared across domains, such as frequency, amplitude, offset, or phase.
+This is particularly valuable in domains with limited data availability and high privacy standards, such as medicine, where our approach offers a promising solution by leveraging readily available data from any domain.
+Overall, our work establishes an essential foundation for future advancements in time series analysis.
 Limitations.
-While OTiS outperforms SOTA models across $10$ tasks, our experiments in low-data regimes suggest that larger pre-training corpora could further enhance its performance.
-Unlike in NLP and CV, where large datasets are curated from web-crawled data, foundational models in time series, including OTiS, still rely on manually curated datasets.
-Future work could explore fully automatic pipelines, e.g.
-using embedding similarity, to filter and rebalance multi-domain time series from the web.
-OTiS could further benefit from processing domain signatures during inference, potentially unlocking zero-shot capabilities, similarly to those seen in foundational models in NLP and CV.
+While OTiS outperforms SoTA models across diverse tasks, scaling studies indicate that larger pre-training corpora could further benefit performance.
+Unlike in NLP and CV, where large corpora are curated from the web, time series foundation models, including OTiS, still rely on manually curated data.
+Future work could explore fully automated pipelines to filter time series from the web and categorise them into domains, e.g.
+using embedding similarities.
+Finally, integrating other modalities, such as imaging or text, is a promising direction for future research.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Large Multi-Domain Pre-Training Corpus
@@
 The full training set includes a class-wise balanced subset (AudioSet-$20$K, $22,176$ clips) and an unbalanced (AudioSet-$2$M $2,042,985$ clips) set.
-For our pre-training corpus, we use the balanced AudioSet-$20$K, which contains $3,491$ mono and $16,123$ stereo recordings, all sampled at $44,100\,$Hz.
+For our pre-training corpus, we use the balanced AudioSet-$20$K, which contains $3,491$ mono and $16,123$ stereo recordings, all sampled at 44.144.144.1\,44.1kHz.
 Electromechanics.
@@
 The Exchange dataset (Lai et al., 2018) records the daily exchange rates of eight different nations, including Australia, Great Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, ranging from $1990$ to $2016$.
-Appendix B Benchmark Dataset Details
-We provide an overview of the datasets used to benchmark our model on $15$ distinct downstream applications in Table 4.
-Our benchmarking experiments include classifcation, regression, and forecasting tasks.
-Table 4: Experimental setup with all conducted downstream applications, evaluation metrics, and datasets used for benchmarking.
+Appendix B Benchmark Details
+To assess the utility of OTiS in real-world settings, we conduct experiments on three key use cases in time series analysis: classification, regression, and forecasting.
+For classification, we perform binary epilepsy detection using EEG (Epilepsy 2001), multi-class fault detection in rolling bearings from vibration signals (FD-B 2016), multi-class hand-gesture classification with accelerometer signals (Gesture 2009), and multi-class muscular disease classification using electromyographie (EMG 2000).
+For regression, we predict five imaging-derived cardiac phenotypes from $12$-lead ECG (LVEDV, LVESV, LVSV, LVEF, LVM 2020).
+For forecasting, we predict electricity transformer temperature (ETT 2021), weather (Weather 2024), and electricity consumption (Electricity 2024).
+We adhere to the established data splitting and evaluation procedures for the classification (Zhang et al., 2022), regression (Turgut et al., 2025), and forecasting (Zhou et al., 2021) tasks.
+We provide an overview of the datasets and baselines used to benchmark our model in Table 4 and Table 5, respectively.
+Table 4: Summary of all datasets used for benchmarking, including evaluation metrics, domains, and dataset details.
 Task Metric Dataset
@@
 Classification ACC EEG Epilepsy 2001 $11,500$ $1$ $178$ 174174174\,174Hz
+TUEV 2016 $112,237$ $91$,00010001,0001 , 000 200200200\,200Hz
 Electromechanics FD-B 2016 $13,640$ $1$ $5,120$ 646464\,64kHz
@@
 Regression $R^{2}$ ECG UK BioBank 2015 $18,926$ $12$ $5,000$ 500500500\,500Hz
-Forecasting MSE Energy ETTh1 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
-ETTh2 2021 $1$ $7$ $17,420$ (hourly) 278278278\,278mHz
+Forecasting MSE Energy ETTh1 2021 $1$ $7$ $17,420$ (hourly) $278\,\mu$Hz
+ETTh2 2021 $1$ $7$ $17,420$ (hourly) $278\,\mu$Hz
 ETTm1 2021 $1$ $7$ $69,680$ (minutely) 1.11.11.1\,1.1mHz
@@
 Weather Weather 2024 $1$ $21$ $52,696$ (minutely) 2.82.82.8\,2.8mHz
-Electricity Electricity 2024 $321$ $1$ $26,304$ (hourly) 278278278\,278mHz
+Electricity Electricity 2024 $321$ $1$ $26,304$ (hourly) $278\,\mu$Hz
+Table 5: Summary of all baseline models used for benchmarking, including pre-training details, domain adaptation methods, and architectural design.
+CL, MDM, and GPT denote contrastive learning, masked data modelling, and generative pre-training, respectively.
+Task Model Pre-training Domain adaptation Architecture
+Method Dataset
+Classification SimCLR 2020 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
+TimesNet 2022a Fine-tuning $2$D-CNN
+CoST 2022 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
+TS2Vec 2022 CL SleepEEG* 2000 Fine-tuning $1$D-CNN
+TF-C 2022 CL SleepEEG* 2000 Fine-tuning Transformer
+Ti-MAE 2023 MDM SleepEEG* 2000 Fine-tuning Transformer
+SimMTM 2024 MDM SleepEEG* 2000 Fine-tuning Transformer
+ST-Transformer 2021 Fine-tuning Transformer
+CNN-Transformer 2022 Fine-tuning CNN w/ Transformer
+FFCL 2022 Fine-tuning CNN w/ LSTM
+SPaRCNet 2023 Fine-tuning $1$D-CNN
+ContraWR 2023 CL Target Fine-tuning Transformer
+BIOT 2024 MDM + Fine-tuning Transformer
+LaBraM 2024 MDM Fine-tuning Transformer
+Regression iTransformer 2023 Fine-tuning Transformer
+ViT 2025 Fine-tuning Transformer
+MAE 2025 MDM Target Fine-tuning Transformer
+CM-AE 2023 MDM and CL Target Fine-tuning $1$D-CNN
+MMCL 2025 MDM and CL Target Fine-tuning Transformer
+Forecasting N-BEATS 2019 Fine-tuning Non-Linear Model
+Autoformer 2021 Fine-tuning Transformer
+TimesNet 2022a Fine-tuning $2$D-CNN
+DLinear 2023 Fine-tuning Linear Model
+PatchTST 2023 MDM Target Fine-tuning Transformer
+Time-LLM 2023 GPT Fine-tuning Transformer
+GPT4TS 2023 GPT Fine-tuning Transformer
+MOMENT 2024 MDM TSP 2024 Fine-tuning Transformer
+MOIRAI 2024 MDM LOTSA 2024 Zero-shot Transformer
+* $371,055$ uni-variate, 2-seconds EEG recordings sampled at a frequency of $100$ Hz.
++ Pre-training corpus spans $6$ EEG datasets with more than $5$ million samples and $13,000$ recording hours
+Pre-training corpus spans $16$ EEG datasets with a total of $2,500$ recording hours
+Model uses Llama-$7$B 2023, pre-trained on $1.4$ trillion text tokens, as encoder.
+Model uses GPT2222 2018, pre-trained on $10$ billion text tokens, as encoder.
+Time Series Pile (TSP) contains $13$ million samples and $1.23$ billion time points from $13$ domains.
+Large-Scale Open Time Series Archive (LOTSA) contains more than $4$ million samples and $27$ billion time points from $9$ domains.
 Appendix C Experiment Details
@@
 We provide the hyperparameters used to pre-train all variants of OTiS in Table 7.
-The hyperparameters used to fine-tune our models for the classification, regression, and forecasting tasks are provided in Table 7, 8, and 9, respectively.
+The hyperparameters used to fine-tune our models for the classification, regression, and forecasting tasks are provided in Table 8, 9, and 10, respectively.
 Table 7: Hyperparameters used for pre-training.
@@
 OTiS-Huge $75$ $32$ $3$e-$4$ $0.1$ $0.75$ $0.2$ $0.1$
-Gesture OTiS-Base $75$ $32$ $1$e-$2$ $0.1$ $0.25$ $0.2$ $0.1$
+Gesture OTiS-Base $75$ $32$ $3$e-$3$ $0.2$ $0.50$ $0.1$ $0.1$
 OTiS-Large $75$ $32$ $3$e-$3$ $0.2$ $0.75$ $0.1$ $0.0$
@@
 Dataset Model Epochs Batch size Base LR NCC $\lambda$ Weight decay
-ETTh1 OTiS-Base $1000$ $1$ $1$e-00 $0.1$ $0.15$
-OTiS-Large $1000$ $1$ $1$e-$1$ $0.2$ $0.15$
-OTiS-Huge $1000$ $1$ $3$e-$1$ $0.1$ $0.15$
-ETTh2 OTiS-Base $1000$ $1$ $1$e-00 $0.2$ $0.25$
-OTiS-Large $1000$ $1$ $1$e-$1$ $0.1$ $0.25$
-OTiS-Huge $1000$ $1$ $3$e-$1$ $0.0$ $0.25$
-ETTm1 OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
-OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
-OTiS-Huge $1000$ $1$ $1$e-$1$ $0.1$ $0.15$
-ETTm2 OTiS-Base $1000$ $1$ $3$e-$1$ $0.1$ $0.25$
-OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.25$
-OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.25$
-Weather OTiS-Base $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
-OTiS-Large $1000$ $1$ $3$e-$1$ $0.2$ $0.15$
-OTiS-Huge $1000$ $1$ $1$e-$1$ $0.2$ $0.05$
-Electricity OTiS-Base $250$ $32$ $3$e-$2$ $0.0$ $0.25$
-OTiS-Large $250$ $32$ $3$e-$2$ $0.0$ $0.15$
-OTiS-Huge $250$ $32$ $3$e-$2$ $0.2$ $0.15$
+ETTh1 OTiS-Base $1000$ $1$ $3$e-$2$ $0.1$ $0.15$
+OTiS-Large $1000$ $1$ $3$e-$3$ $0.2$ $0.15$
+OTiS-Huge $1000$ $1$ $1$e-$2$ $0.1$ $0.15$
+ETTh2 OTiS-Base $1000$ $1$ $3$e-$2$ $0.2$ $0.25$
+OTiS-Large $1000$ $1$ $3$e-$3$ $0.1$ $0.25$
+OTiS-Huge $1000$ $1$ $1$e-$2$ $0.0$ $0.25$
+ETTm1 OTiS-Base $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
+OTiS-Large $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
+OTiS-Huge $1000$ $1$ $3$e-$3$ $0.1$ $0.15$
+ETTm2 OTiS-Base $1000$ $1$ $1$e-$2$ $0.1$ $0.25$
+OTiS-Large $1000$ $1$ $3$e-$3$ $0.2$ $0.25$
+OTiS-Huge $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
+Weather OTiS-Base $1000$ $1$ $1$e-$2$ $0.2$ $0.25$
+OTiS-Large $1000$ $1$ $1$e-$2$ $0.2$ $0.15$
+OTiS-Huge $1000$ $1$ $3$e-$3$ $0.2$ $0.05$
+Electricity OTiS-Base $250$ $32$ $1$e-$3$ $0.0$ $0.25$
+OTiS-Large $250$ $32$ $1$e-$3$ $0.0$ $0.15$
+OTiS-Huge $250$ $32$ $1$e-$3$ $0.2$ $0.15$
 Appendix D Computation Costs
@@
 Appendix E Domain Signature Analysis
-To analyse the domain signatures, we reduce the dimensionality of the domain-specific variate embeddings by employing a principal component analysis (PCA).
-Our analysis shows that OTiS unifies time series from diverse domains into a meaningful latent space, while accurately capturing the inter-variate relationships within a domain.
+We analyse the domain signatures, i.e.
+the domain-specific variate embeddings, by employing a principal component analysis (PCA) and embedding similarity analysis.
+Our analyses show that OTiS unifies time series from diverse domains into a meaningful latent space (see Section E.1), while accurately capturing the inter-variate relationships (see Section E.2.1) and temporal patterns within a domain (see Section E.2.2).
 E.1 Inter-Domain Analysis
 A visualisation of all domain-specific variate embeddings learned during pre-training is provided in Figure 6.
-We find that OTiS learns a meaningful latent space, where embeddings of domains with shared high-level semantics cluster closely together.
+OTiS learns a meaningful latent space, where embeddings of domains with shared high-level semantics cluster closely together.
 (a) $3$D projection.
@@
 E.2 Intra-Domain Analysis
-A visualisation of the domain-specific variate embeddings for EEG and ECG are given in Figure 7 and Figure 8, respectively.
-We find that OTiS preserves low-level semantics specific to each domain, accurately capturing the relationships between variates.
-(a) True electrode layout.
-(b) Implicitly learned electrode layout.
-Figure 7: $3$D PCA projections of the variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training.
-The labels of each node correspond to the electrode names.
+E.2.1 Inter-Variate Relationships
+EEG-Specific Variate Embeddings.
+In medicine, an electroencephalography (EEG) is used to record brain activity via electrodes attached to the scalp.
+The spatial arrangement of these electrodes is defined by established protocols, such as the international $10$-$20$ system for EEG recordings (Homan et al., 1987).
+Hence, in an EEG recording, each variate corresponds to the brain activity captured by a single electrode with a well-defined spatial position.
+We hypothesise that the spatial arrangement of these electrodes is recoverable from the EEG-specific variate embeddings, assuming that recordings from electrodes in close proximity exhibit higher correlations than those from electrodes placed farther apart.
+To test this hypothesis, we analyse the latent space of EEG-specific variate embeddings learned during pre-training and fine-tuning, as described in the following.
+(a) $3$D spatial arrangement of EEG electrodes defined by the international $10$-$20$ system for EEG recordings.
+(b) First three principal components of EEG-specific variate embeddings.
+Figure 7: Analysis of the variate embeddings for $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training. The label of each coordinate corresponds to the electrode name.
+Assume 3D electrode coordinates of the international $10$-$20$ system for EEG recordings (Homan et al., 1987) to be defined in Euclidean space $\mathbb{E}_{Y}^{3}$ (see Figure 7(a)).
+We first determine the first three principal components of the EEG-specific variate embeddings, thus projecting them into a Euclidean space $\mathbb{E}_{X}^{3}$ (see Figure 7(b)).
+Then, we perform a multivariate linear regression
+$\bm{Y}=\bm{1}\beta_{0}+\bm{X}\bm{B}+\bm{\epsilon}\in\mathbb{R}^{N\times 3}\quad\text{with}\quad\beta_{0}\in\mathbb{R}^{1\times 3},\bm{X}\in\mathbb{R}^{N\times 3},\bm{B}\in\mathbb{R}^{3\times 3},\bm{\epsilon}\in\mathbb{R}^{N\times 3},$ (8)
+where $\bm{1}\in\mathbb{R}^{N\times 1}$ is a vector of ones and $N$ denotes the number of electrodes, to align the first three principal components in $\mathbb{E}_{X}^{3}$ (here, $\bm{X}$) with the $3$D electrode coordinates in $\mathbb{E}_{Y}^{3}$ (here, $\bm{Y}$).
+Finally, to quantify this very alignment, we determine the coefficient of determination $R^{2}\in[0,1]$.
+Note that $R^{2}=1$ represents a perfect alignment, i.e.
+$\mathbb{E}_{Y}^{3}=\mathbb{E}_{X}^{3}$, where the first three principal components of the EEG-specific variate embeddings only need to be shifted and scaled to recover the spatial arrangement of the electrodes (i.e.
+$\epsilon$ is a zero matrix).
+We analyse (i) variate embeddings of $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training and (ii) variate embeddings of previously unseen $10$-$20$ system EEG recordings with $32$ electrodes learned during fine-tuning.
+We find that the first three principal components explain $74.7\,\%$ of the variance in (i) and $87.9\,\%$ in (ii).
+Moreover, we observe $R^{2}$ values of $0.81$ for (i) and $0.95$ for (ii), indicating an accurate alignment of EEG-specific variate embeddings with the electrode layout.
+A $2$D visualisation of this alignment is provided in Figure 8.
+Overall, these findings confirm our initial hypothesis that the spatial arrangement of EEG electrodes can be precisely recovered from EEG-specific variate embeddings.
+(a) Variate embeddings of $10$-$20$ system EEG recordings with $19$ electrodes learned during pre-training.
+(b) Variate embeddings of previously unseen $10$-$20$ system EEG recordings with $32$ electrodes learned during fine-tuning.
+Figure 8: First two principal components of EEG-specific variate embeddings, overlaid on the true EEG electrode layout.
+The embeddings accurately reflect the spatial arrangement of the EEG electrodes, confirmed by high correlations (R2) between the principal components $\bullet$ and the electrode arrangement $\circ$.
+ECG-Specific Variate Embeddings.
+Another application in medicine is electrocardiography (ECG), which is widely used to measure the electrical activity of the heart via electrodes placed on the human torso.
+The standard $12$-lead ECG is composed of $3$ limb leads (leads I, II, and III), $3$ augmented limb leads (leads aVR, aVL, and aVF), and $6$ precordial leads (leads V1 to V6).
+The limb leads (or Einthoven leads Einthoven et al.
+(1902)) reflect the potential differences between pairs of electrodes attached to the limbs (right arm: RA, left arm: LA, and left leg: LL): I, II, and III represent the potential difference between RA and LA, RA and LL, and LA and LL, respectively.
+Together, these limb leads form the so called Einthoven triangle Kligfield et al.
+(2007), spanned by the three vectors RA $\rightarrow$ LA, RA $\rightarrow$ LL, and LA $\rightarrow$ LL.
+Similarly, the augmented limb leads (or Goldberger leads Goldberger (1942)) also reflect potential differences between pairs of electrodes attached to the limbs.
+In contrast to the previous leads, however, the precordial leads (or Wilson leads Wilson et al.
+(1934)) are directly measured from the electrodes placed on the chest.
+When analysing the latent space of the ECG-specific variate embeddings learned during pre-training, we can clearly observe these relationships among the $12$ leads, as depicted in Figure 9.
+For instance, the limb leads I, II, and III form a planar $2$D triangle, thus successfully reflecting Einthoven’s triangle Kligfield et al.
+(2007).
+Additionally, the precordial leads V1111 to V6666, arranged on the rib cage from the sternum to the mid-axillary line, represent the $3$D view of the human heart.
 (a) Side view.
@@
 (c) Top view.
-Figure 8: $3$D PCA projections of the variate embeddings for standard $12$-lead ECG learned during pre-training.
-The embeddings accurately reflect the true physiological structure of the ECG leads.
+Figure 9: Principal component analysis of the variate embeddings for standard $12$-lead ECG learned during pre-training.
+Their first three components, shown in (a), (b), and (c), accurately reflect the true physiological structure of ECG leads.
 The V1111-V6666 leads, arranged on the rib cage from the sternum to the mid-axillary line, represent a $3$D view of the human heart.
-In contrast, the I-II-III leads and aVR-aVL-aVF leads, derived from electrodes placed on one foot and both arms, form a planar $2$D triangle.
-Appendix F Forecast Visualisation
+The I-II-III leads and aVR-aVL-aVF leads, derived from electrodes placed on one foot and both arms, form a planar $2$D triangle.
+Weather-Specific Variate Embeddings.
+We focus the following analysis on the Weather-specific variate embeddings learned during fine-tuning.
+The central question is whether OTiS can learn domain-specific knowledge - in this case, for the weather domain - from limited data seen only during fine-tuning.
+To investigate this, we compute the cosine similarity for all pairs of Weather-specific variate embeddings, as summarised in Figure 10.
+Note that these embeddings were randomly initialised and learned specifically for the Weather 2024 dataset during the forecasting task described in Section B.
+The Weather variates span diverse climatological categories, including temperature (T, Tpot, Tdew, Tlog), humidity (rh, VPmax, VPact, VPdef, sh, H2OC), wind (wv, max.
+wv, wd), radiation (SWDR, PAR, max.
+PAR), pressure (p, rho), and precipitation (rain, raining).
+Our analysis demonstrates that OTiS effectively captures complex relationships among these distinct climatological indicators, as detailed in the following discussion.
+High positive similarities typically indicate relationships within a single climatological category.
+For example, we observe strong similarities among temperature variates, humidity variates, radiation variates, pressure variates, and precipitation variates.
+These results are expected, as these variates all describe different aspects of the same category and often fluctuate together.
+Additionally, subtle variations in the similarity scores reveal how, for instance, dew point temperature (Tdew) depends not only on temperature but also on other factors, such as humidity (rh).
+High negative similarities typically represent relationships across climatological categories.
+For example, consider the inverse relationship between vapor pressure deficit (VPdef) and relative humidity (rh), defined as:
+$\text{VPdef}=\text{SVP}\left(1-\frac{\text{rh}}{100}\right),$ (9)
+where SVP [mBar] denotes the saturation vapor pressure.
+Our analysis showcases that OTiS correctly captures this negative correlation, as well as other relationships across categories.
+These include the inverse correlation between strong winds (max.
+wv) and low air pressure (p), and between extended precipitation (raining) and lower incoming radiation (SWDR).
+Figure 10: Cosine similarity matrix of Weather-specific variate embeddings.
+The ordering of the variates was modified for visualisation purposes.
+Areas with high positive and high negative similarity are exemplary framed in yellow.
+OTiS successfully learns the non-linear relationships between climatological indicators of the Weather 2024 dataset.
+E.2.2 Temporal Patterns
+The central question is whether OTiS learns only the correlations across variates or also captures temporal patterns within a variate.
+To investigate this, we conduct forecasting experiments on uni-variate sine waves with distinct frequencies ranging from 222\,2Hz to 100100100\,100Hz.
+In this uni-variate setting, we ensure that our model does not leverage information from other variates.
+For these experiments, OTiS remains frozen after pre-training, with only the randomly initialised domain-specific variate embedding (a single embedding for uni-variate data, totalling less than 0.20.20.2\,0.2k trainable parameters) being fine-tuned on 505050\,50Hz sine waves.
+During inference, we forecast unseen sine waves with frequencies of 222\,2Hz, 282828\,28Hz, 606060\,60Hz, and 100100100\,100Hz, using the domain-specific variate embedding learned on 505050\,50Hz sine waves.
+As depicted in Figure 11, the experiments reveal that OTiS captures not only inter-variate relationships, but also temporal patterns that generalise to unseen time series.
+Figure 11: Visualisation of OTiS-Base forecast predictions on sine waves.
+A forecasting horizon of $96$ time points is predicted from the past $336$ time points.
+Ground truth in blue, prediction in orange.
+Areas highlighted in grey are not visible to the model.
+During fine-tuning, OTiS remains frozen after pre-training, with only the randomly initialised domain-specific variate embedding ($<0.2\,$k parameters) being trained.
+During inference, OTiS is fully frozen and forecasts are made on unseen sine waves.
+The domain-specific variate embedding trained on 505050\,50Hz sine waves effectively captures temporal patterns that generalise to sine waves of previously unseen frequencies.
+Appendix F Zero-Shot Capabilities
+To analse the quality of the time series features extracted by OTiS, we conduct zero-shot classification experiments using the Epilepsy 2001, FD-B 2016, Gesture 2009, and EMG 2000 datasets.
+These datasets ensure high versatility in our analysis, as they differ notably in the number of variates, time points, sampling frequency, as well as number of classes, and cover distinct domains and applications.
+For these zero-shot experiments, OTiS is evaluated directly after pre-training without any fine-tuning, utilising randomly initialised variate embeddings.
+As no classification head is employed, the encoder’s output tokens are averaged to obtain a global representation for each sample.
+To generate class logits, pairwise cosine similarity is computed between the global representation of a test sample and each class representation, i.e.
+the mean of all global representations of the training samples from that class.
+The class with the highest logit is assigned to the test sample.
+As illustrated in Figure 12, our analysis demonstrates that OTiS extracts distinct representations for different classes, even without domain-specific fine-tuning.
+This capability translates to zero-shot classification accuracies of $93.70\,\%$ for Epilepsy, $57.87\,\%$ for FD-B, $51.67\,\%$ for Gesture, and $95.12\,\%$ for EMG.
+Comparing these results with those from fine-tuning in Figure 13, we observe only slight performance decreases for Epilepsy ($-0.55\,\%$) and EMG ($-2.44\,\%$), but more substantial gaps for FD-B ($-41.37\,\%$) and Gesture ($-11.94\,\%$).
+A closer comparison of the latent space for FD-B between the zero-shot setup (Figure 12(b)) and the fine-tuning setup (Figure 13(b)) reveals that the performance gap arises due to a partial overlap of inputs from classes $1$ and $2$ in the zero-shot setup.
+Similarly, the zero-shot latent space shows poor clustering for the eight classes in Gesture (Figure 12(c)), with notable improvements for classes 00 and $7$ after fine-tuning (Figure 13(c)).
+Furthermore, since our pre-training corpus includes EEG (TDBrain 2022 and SEED 2015) and Electromechanics (FD-A 2016) data, we investigate whether performance gains can be achieved by using domain-specific variate embeddings, learned during pre-training, instead of randomly initialised ones.
+As expected, leveraging these domain-specific embeddings improve the quality of the extracted time series features, resulting in zero-shot classification accuracies of $95.18\,\%$ for Epilepsy ($+1.48\,\%$) and $61.32\,\%$ for FD-B ($+3.45\,\%$).
+Interestingly, the improved zero-shot performance for Epilepsy even surpasses that of fine-tuning.
+Overall, these zero-shot experiments indicate a high quality of the readily available time series features extracted by OTiS, which can be further refined through domain-specific knowledge, including domain-specific variate embeddings or fine-tuning.
+(a) Epilepsy.
+(b) FD-B.
+(c) Gesture.
+(d) EMG.
+Figure 12: First two principal components of zero-shot representations generated by OTiS-Base.
+In this setup, OTiS is evaluated after pre-training without any fine-tuning, utilising randomly initialised variate embeddings.
+As no classification head is employed, the output tokens of the encoder are averaged to obtain a global representation.
+OTiS generates distinct representations for different inputs, even across domains and tasks, highlighting the quality of its extracted time series features.
+(a) Epilepsy.
+(b) FD-B.
+(c) Gesture.
+(d) EMG.
+Figure 13: First two principal components of fine-tuned representations generated by OTiS-Base.
+In this setup, OTiS is fully fine-tuned, along with a classification head, after pre-training.
+To enable a fair comparison, the global representations obtained by averaging the encoder’s output tokens are visualised.
+Incorporating domain-specific knowledge through fine-tuning enables OTiS to effectively tailor its feature extraction to the specific domain and task.
+Appendix G Additional Ablation Studies
+G.1 Dual Masking Strategy
+To enhance OTiS’ foundational capabilities for general time series analysis, we utilise a dual masking strategy during pre-training, as described in Section 3.4.
+Specifically, we alternate between two masking schemes during pre-training: random masking (randomly masking across variate and temporal dimension) and post-fix masking (masking the second half of the temporal dimension).
+To determine the optimal balance between these two schemes, we evaluate different compositions of the dual masking strategy across the general use cases in time series analysis.
+Our analysis reveals that a combination of $75\,\%$ random masking and $25\,\%$ post-fix masking consistently yields the best downstream performance across all tasks, as illustrated in Figure 14.
+Figure 14: Ablation study on the composition of the dual masking strategy.
+Error bars represent the standard deviation across $5$ seeds.
+A combination of $75\,\%$ random masking and $25\,\%$ post-fix masking during pre-training consistently yields the best downstream performance across tasks.
+G.2 Pre-Training Strategy
+To explore whether domain-specific pre-training offers advantages over pre-training on diverse time series across domains, we analyse various training strategies for EEG event type classification using the TUEV 2016 dataset, as summarised in Table 12.
+In particular, we evaluate three different training strategies for our model: fully supervised training from scratch (OTiS-Basew/o ${}_{\text{w/o pre-training}}$), pre-training exclusively on ${}_{\text{EEG}}$), and pre-training across domains (OTiS-Base).
+Additionally, we analyse whether domain-specific variate embeddings learned during pre-training are required for competitive downstream performance, or if randomly initialised variate embeddings (rVE) are sufficient.
+We compare OTiS against both task-specific and foundation models that are particularly designed for EEG analysis.
+The task-specific models (either fully supervised or pre-trained and fine-tuned on the target dataset) include ST-Transformer (Song et al., 2021), CNN-Transformer (Peh et al., 2022), FFCL (Li et al., 2022), SPaRCNet (Jing et al., 2023), and ContraWR (Yang et al., 2023).
+The foundation models (pre-trained on a large medical time series corpus and fine-tuned on the target dataset) include BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024).
+Further details on training strategies and architectural design are provided in Table 5 of Section B.
+The experiments show that pre-trained models consistently outperform fully supervised models trained from scratch.
+However, our fully supervised model, OTiS-Basew/o ${}_{\text{w/o pre-training}}$, is superior to task-specific models and performs on par with the foundational BIOT model, indicating an efficient interplay between the domain-specific tokeniser and the encoder.
+We also observe that competitive performance can be achieved even without explicit domain knowledge, as indicated by the results of OTiS-BaseEEG w/ ${}_{\text{EEG w/ rVE}}$.
+Interestingly, we find that pre-training OTiS exclusively on EEG data does not yield better results compared to pre-training across diverse domains.
+These observations emphasise that expanding the pre-training corpus with time series from diverse domains enhances the overall quality of extracted time series features, which is especially beneficial when domain-specific data is scarce.
+Optimal performance within a specific domain requires pre-training on large, domain-specific corpora, which provide comprehensive domain knowledge and translate into superior downstream results, as evidenced by the performance of LaBraM.
+However, this approach is often hindered by the limited availability of domain-specific data, particularly in medical applications.
+In such cases, our approach offers a promising solution to address data scarcity by leveraging readily available data from other domains.
+Table 12: Ablation study on training strategies for EEG event type classification on TUEV 2016.
+Mean and standard deviation is reported across $5$ seeds.
+Best score in bold, second best underlined.
+Note that PT refers to pre-trained models.
+Expanding the EEG pre-training corpus of OTiS with time series from other domains improves the overall quality of the extracted EEG features.
+Methods PT Params Balanced ACC $\uparrow$ Cohen’s Kappa $\uparrow$ Weighted F1 $\uparrow$
+ST-Transformer 2021 3.53.53.5\,3.5M 0.3984 $\pm$ 0.0228 0.3765 $\pm$ 0.0306 0.6823 $\pm$ 0.0190
+CNN-Transformer 2022 3.23.23.2\,3.2M 0.4087 $\pm$ 0.0161 0.3815 $\pm$ 0.0134 0.6854 $\pm$ 0.0293
+FFCL 2022 2.42.42.4\,2.4M 0.3979 $\pm$ 0.0104 0.3732 $\pm$ 0.0188 0.6783 $\pm$ 0.0120
+SPaRCNet 2023 0.80.80.8\,0.8M 0.4161 $\pm$ 0.0262 0.4233 $\pm$ 0.0181 0.7024 $\pm$ 0.0104
+ContraWR 2023 1.61.61.6\,1.6M 0.4384 $\pm$ 0.0349 0.3912 $\pm$ 0.0237 0.6893 $\pm$ 0.0136
+BIOT+ 2024 3.23.23.2\,3.2M 0.5281 $\pm$ 0.0225 0.5273 $\pm$ 0.0249 0.7492 $\pm$ 0.0082
+LaBraM 2024 369369369\,369M 0.6616 $\pm$ 0.0170 0.6745 $\pm$ 0.0195 0.8329 $\pm$ 0.0086
+OTiS-Basew/o ${}_{\text{w/o pre-training}}$* 888\,8M 0.5361 $\pm$ 0.0350 0.5183 $\pm$ 0.0316 0.7642 $\pm$ 0.0157
+[…]
+OTiS-BaseEEG w/ ${}_{\text{EEG w/ rVE}}$ $8$ M 0.5413 $\pm$ 0.0302 0.5631 $\pm$ 0.0299 0.7860 $\pm$ 0.0120
+OTiS-Base 888\,8M 0.5743 $\pm$ 0.0257 0.5913 $\pm$ 0.0146 0.8004 $\pm$ 0.0071
+OTiS-Basew/ ${}_{\text{w/ rVE}}$ $8$ M 0.5728 $\pm$ 0.0134 0.5772 $\pm$ 0.0281 0.7922 $\pm$ 0.0127
++ Model was pre-trained on $6$ EEG datasets, including TUEV 2016, with more than $13,000$ recording hours.
+Model was pre-trained on $16$ EEG datasets, including data from the TUH EEG corpus 2016, with more than $2,500$ recording hours.
+* Model was randomly initialised and trained fully supervised.
+Model was pre-trained on EEG datasets of our pre-training corpus, i.e. TDBrain 2022 and SEED 2015, totalling $125$ recording hours.
+Variate embeddings (VE) are randomly (r) initialised before for fine-tuning.
+Appendix H Forecast Visualisation
 We visualise the performance of our model on $6$ forecasting benchmarks in Figure 15.